### v2

# Fourier Analysis of Finite Difference Schemes for the Helmholtz Equation in 1D with Dirichlet Conditions: Sharp Estimates and Relative Errorsthanks: 2020 Mathematics Subject Classification: 65N06, 65N12, 65N15

Abstract
We propose an approach based on Fourier analysis to wavenumber explicit sharp estimation of absolute and relative errors of finite difference methods for the Helmholtz equation in 1D with Dirichlet boundary conditions and general source terms. We use the approach to analyze the classical centred scheme. For the Fourier interpolants of the discrete solution with homogeneous (or inhomogeneous) Dirichlet conditions, we show rigorously that the worst case attainable convergence order of the absolute error is $k^{2}h^{2}$ (or $k^{3}h^{2}$) in the $L^{2}$-norm and $k^{3}h^{2}$ (or $k^{4}h^{2}$) in the $H^{1}$-semi-norm, and that of the relative error is $k^{3}h^{2}$ in both $L^{2}$- and $H^{1}$-semi-norms. Even though the classical centred scheme is well-known, it is the first time that such sharp estimates of absolute and relative errors are obtained. We show also that the Fourier analysis approach can be used as a convenient visual tool for evaluating finite difference schemes in presence of source terms, which is beyond the scope of dispersion analysis.
1 Introduction
The Helmholtz equation $\Delta u+k^{2}u=f$ is a common model of time-harmonic waves in acoustics [24], geophysics[37] and electrical engineering[4], etc. Standard discretization of the Helmholtz equation suffers from the so-called “pollution effect”, explained by [11] as “the accuracy of the numerical solution deteriorates with increasing non-dimensional wave number kk”. It is proved in [26] that the linear finite element method for the 1D problem on $(0,1)$ with $u(0)=0$ and $u^{\prime}(1)-\mathrm{i}ku(1)=0$ has the error in the $H^{1}$-norm bounded from above by $C(hk+h^{2}k^{3})\|f\|_{L^{2}}$ with $C>0$ independent of $h$, $k$ and $f$, where $Ch^{2}k^{3}\|f\|_{L^{2}}$ is called the “pollution term” which indicates the accuracy deterioration with the “rule of thumb” keeping $kh$ constant. Since then, a lot of progress has been made in wavenumber explicit error estimates of finite element methods [34, 46, 51, 12, 32, 39], including the recent extension to heterogeneous media [6, 8, 21, 30] and multiscale methods [36, 18, 22, 7, 33].
While finite element methods are more advanced and powerful, finite difference methods are still used for their simplicity in some applications [40, 23, 2, 41, 35]. In particular, it was found in [41] for full waveform inversion that “a 27-point stencil on a regular Cartesian grid” is “more computationally efficient” than “a $P_{3}$ finite element method on $h$-adaptive tetrahedral mesh”. It is pressing to have a theoretical support for methods in use. But wavenumber explicit error estimates are rarely seen in the literature of finite difference methods [38, 43, 16, 45, 48]. Error estimates of some compact fourth order schemes in 2D are given in [19] based on the assumption that the coefficient matrix of the discretized system is positive definite (which is true only for small wavenumber $k$). In [44], arbitrarily high order and dispersion free schemes in 1D are derived, and an error analysis following the approach of [19] is given. Recently, in [9] it is shown for a dispersion free 3-point scheme in 1D that the maximum-norm error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor. An earlier dispersion free scheme in 1D was proposed in the seminal paper [5] and analyzed in the finite element framework giving the upper bound of the $H^{1}$-semi-norm error as $ch\|f\|_{H^{1}}$ for the scattering problem (transparent on one end and Dirichlet on the other) with $c$ seemingly $O(k)$.
It has been well noted that the pollution effect is intuitively related to the fact that the numerical solution “has a wavelength that is different from the exact one”, called “dispersion[11]. Therefore, [11] advocated “a tool in order to measure quantitatively the dispersion for various of the proposed methods as a measure of the pollution”. “This measure allows to compare the different methods that have been proposed to reduce the pollution and compare their efficiency.” The tool called “dispersion analysis[1] had been used before mainly for time-dependent problems [42].
The idea of dispersion analysis is to find Fourier modes e.g. $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}$ satisfying the discretized homogeneous Helmholtz equation $\Delta^{h}u^{h}+k^{2}u^{h}=0$ and from the resulting constraint on $\boldsymbol{\xi}$ find the discrete wavenumber $k^{h}:=|\boldsymbol{\xi}|$ as a function of the direction $\mathbf{l}:=\boldsymbol{\xi}/|\boldsymbol{\xi}|$. For example, if the original Helmholtz equation is used instead of the discretized one, then we find the constraint $|\boldsymbol{\xi}|=k$ defining a spherical surface in the space of Fourier frequencies $\boldsymbol{\xi}$. Such Fourier modes $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}=\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ are called “plane waves”. The “phase difference” $|k-k^{h}|$ as a function of the unit direction $\mathbf{l}$ is a measure of the dispersion error between the original plane wave $\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ and the discretized dispersive ($k^{h}\neq k$) plane wave $\mathrm{e}^{\mathrm{i}k^{h}\mathbf{l}\cdot\mathbf{x}}$.
It is shown in [27] for polynomial finite element methods in 1D that the pollution term is of the same order as the phase difference. That partially explains why reducing the phase difference leads to reducing the pollution effect. Sharp dispersion error estimates are obtained in [3]. By providing a practical criterion, dispersion analysis can also guide optimization of the discretization parameters; see [50] for the continuous interior penalty finite element method and [40, 47, 10, 9] for finite difference methods. Recently, dispersion reduction has been generalized in [31] to unstructured meshes by minimizing the residual of the original plane waves in the discretized equation; see also the earlier work [17]. A similar approach based on minimizing “average truncation error of plane waves” is adopted in [16] for some high order finite difference methods.
Albeit intuitive, it is not trivial to translate the “dispersion error” into the “pollution effect”. As remarked in [25]:“dispersion analysis gives valuable information on several physical phenomena inherent to the discrete solution and thus leads to qualitative insight into the sources of numerical error,” but “it does not yield, by its nature, quantitative statements on the numerical error itself”. That is why so many efforts have been made in wavenumber explicit error estimation, as we mentioned above.
Moreover, dispersion analysis is insufficient because it considers only the zero source problem and the Fourier modes constrained by the dispersion relation, while nonzero source terms and the other Fourier modes stimulated by the source term are not taken into account. We shall see in Section 4 that three dispersion free schemes in 1D can behave differently with source term. Indeed, in [5] a discretization in 1D was proposed with zero truncation error “for as many right hand sides as possible”, not only the zero right hand side. An eigenvalue (Fourier) analysis is carried out in [13] for the Dirichlet boundary value problem with a point source.
Inspired by [13] which calculated the eigenvalue difference of the continuous and discrete Helmholtz operators, we develop here an approach to wavenumber explicit error estimation based on Fourier analysis for general smooth source terms and Dirichlet boundary conditions in 1D. We analyze in detail a classical centred scheme in 1D because it is simple yet not fully understood. Actually, we have not found any other wavenumber explicit estimate for finite difference schemes than that in [9], and in particular none for the classical centred scheme, though the folklore says a $p$th order scheme has the (relative) error $O(k^{p+1}h^{p})$ (see e.g. [13]). So we believe our main results (see the Abstract for a preview) are novel.
In particular, all the upper bounds of our error estimates are attainable and the equality cases are given explicitly. The pollution term is located in Fourier frequencies, and its exact order in $k$ and $h$ is found by two-sided bounds. To the best of our knowledge, it is the first time that such wavenumber explicit error estimates of such sharpness with source terms are obtained for finite difference methods.
We also give a sharp estimate of relative errors with source term. Wavenumber explicit a priori estimates of relative errors have been much less studied than of absolute errors. Recent progress for finite element methods can be found in [29, 20].
Moreover, by plotting the symbol errors we can see clearly how the error is distributed over Fourier frequencies. In this sense, it becomes a practical tool for comparing different schemes, e.g. those of the same order and those before and after dispersion correction, which complements the dispersion analysis for working with or without source term. This aspect has been briefly demonstrated in [49] for some compact high order schemes, where sharp orders in $k$, $h$ are found in Fourier domain and verified numerically.
In the following sections, we first introduce the model problem and Fourier decomposition in Section 2, then carry out the theoretical analysis of the classical scheme in Section 3, and illustrate the practical aspect of the tool in Section 4. Some discussions are included at the end.
2 Model problem in 1D and downsampling error
We first introduce the 1D Helmholtz equation in a closed cavity, and the Fourier form of the solution. Then we analyze the component of the discretization error due to sampling of the solution on grids.
2.1 Model problem
Let $k>0$ be the wavenumber. The Helmholtz equation with Dirichlet boundary conditions in 1D is
$u^{\prime\prime}+k^{2}u=f\text{ in }(0,1),\quad u(0)=g_{0},\quad u(1)=g_{1}.$ (1)
For well-posedness of (1), we assume that $k/\pi$ is not an integer. By linearity, we separate the two cases:
  • (i)
    homogeneous Dirichlet boundary condition $g_{0}=g_{1}=0$, and
  • (ii)
    zero source $f=0$.
In case (i) $g_{0}=g_{1}=0$, we assume $f\in H_{0}^{p}(0,1)$ with $p\geq 1$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. In principle, $f\in L^{2}(0,1)$ is sufficient to guarantee that $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. The smoother source $f$ in $H_{0}^{p}$ or even $C_{c}^{\infty}$ can be thought as an arbitrarily close approximation of the general $L^{2}$ source. For the corresponding theory, we refer to [15]. Smoothness of $f$ also facilitates analysis of discretization schemes for (1). For example, [5] used the condition $f\in H^{1}(0,1)$. Here $f\in H_{0}^{p}(0,1)$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$ admit odd extension to the domain (1,1)(-1,1) and the convergent sine series in the corresponding spaces $H^{p}(-1,1)$ and $H^{2}(-1,1)$:
$f(x)=\sum_{(\xi/\pi)=1}^{\infty}\hat{f}(\xi)\sin(\xi x),\quad\hat{f}(\xi)=2\int_{0}^{1}f(x)\sin(\xi x)~\mathrm{d}x,$
$u(x)=\sum_{(\xi/\pi)=1}^{\infty}\hat{u}(\xi)\sin(\xi x),\quad\hat{u}(\xi)=2\int_{0}^{1}u(x)\sin(\xi x)~\mathrm{d}x.$
From (1) and the sine series, we have $(k^{2}-\xi^{2})\hat{u}(\xi)=\hat{f}(\xi)$. Well-posedness of (1) amounts to $\lambda(\xi):=k^{2}-\xi^{2}\neq 0$ for all $\xi\in\pi\mathbb{N}$. In case (ii) $f=0$, the exact solution is
$u(x)=\frac{g_{1}-g_{0}\cos k}{\sin k}\sin(kx)+g_{0}\cos(kx).$ (2)
To analyze the discretization error for the homogeneous problem (ii), one just needs to find the closed form solution of the discretized problem. Usually, the discretization incurs a perturbation of the wavenumber $k$ in the solution basis $\{\sin(kx),\cos(kx)\}$, which explains the error, and requires dispersion analysis.
2.2 Downsampling error
We shall discretize (1) on uniform grids with grid points $x_{j}:=jh$, $j=0,\ldots,N$ where $Nh=1$. In a finite difference scheme, $f$ is simply evaluated at the grid points $x_{j}$ to give $f^{h}(x_{j}):=f(x_{j})$. This causes aliasing of the sinusoidal modes in case (i), and gives the discrete sine transform
$f^{h}(x_{j})=\sum_{(\xi/\pi)=1}^{N-1}\widehat{f^{h}}(\xi)\sin(\xi x_{j}),\quad\widehat{f^{h}}(\xi)=2\sum_{j=1}^{N-1}h\sin(\xi x_{j})f^{h}(x_{j})=\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi),$
where the last equality comes from the fact $f^{h}(x_{j})=f(x_{j})$ and the rearrangement of
$f(x_{j})=\sum_{(\xi/\pi)=1}^{\infty}\hat{f}(\xi)\sin(\xi x_{j})=\sum_{(\xi/\pi)=1}^{N-1}\left(\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\right)\sin(\xi x_{j})$
permitted by the absolute convergence
$\sum_{(\xi/\pi)=1}^{\infty}|\hat{f}(\xi)|\leq\sqrt{\sum_{(\xi/\pi)=1}^{\infty}\frac{1}{\xi^{2}}}\cdot\sqrt{\sum_{(\xi/\pi)=1}^{\infty}|\xi\hat{f}(\xi)|^{2}}<\infty.$
Indeed, it is well known that $\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\frac{\pi^{2}}{6},$ and by Parseval’s identity $\sum_{(\xi/\pi)=1}^{\infty}|\xi\hat{f}(\xi)|^{2}=2\int_{0}^{1}|f^{\prime}|^{2}~\mathrm{d}x.$
Let $u^{h}$ be the finite difference solution defined at the grid points. To calculate the error $e^{h}:=u-u^{h}$ for case (i), we extend $u^{h}$ to the domain $(0,1)$ by the discrete sine transform [^1] The sine interpolation is chosen to facilitate our analysis because the continuous solution $u$ is expanded to sine series. In [5], piecewise linear interpolation is used for analysis of a linear finite element method.
$u^{h}(x)=\sum_{(\xi/\pi)=1}^{N-1}\widehat{u^{h}}(\xi)\sin(\xi x),\quad x\in(0,1),\quad\widehat{u^{h}}(\xi):=2\sum_{j=1}^{N-1}h\sin(\xi x_{j})u^{h}(x_{j}).$
So the error for case (i) can be expanded under the sine basis into
$e^{h}(x)=\sum_{(\xi/\pi)=1}^{N-1}(\hat{u}-\widehat{u^{h}})(\xi)\sin(\xi x)+\sum_{(\xi/\pi)=N}^{\infty}\hat{u}(\xi)\sin(\xi x)=:e_{1}^{h}(x)+e_{2}^{h}(x).$ (3)
We first estimate the $H^{1}$-semi-norm of $e_{2}^{h}$. Note that
$(e_{2}^{h})^{\prime}(x)=\sum_{(\xi/\pi)=N}^{\infty}\hat{u}(\xi)\xi\cos(\xi x)=\sum_{(\xi/\pi)=N}^{\infty}\frac{\xi}{k^{2}-\xi^{2}}\hat{f}(\xi)\cos(\xi x).$ (4)
Parseval’s identity for the above cosine series reads
$|e_{2}^{h}|_{1}^{2}:=\int_{0}^{1}|(e_{2}^{h})^{\prime}|^{2}~\mathrm{d}x=\frac{1}{2}\sum_{(\xi/\pi)=N}^{\infty}\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}|\hat{f}(\xi)|^{2}.$ (5)
Suppose $\mu:=\frac{kh}{2}\leq C_{\mu}$ for a constant $C_{\mu}\in(0,\frac{\pi}{2})$. Then $k<N\pi$, and for $\xi\geq N\pi$,
$\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}\leq\frac{(N\pi)^{2}}{((N\pi)^{2}-k^{2})^{2}}=\frac{\pi^{2}h^{2}}{(\pi^{2}-k^{2}h^{2})^{2}}\leq\frac{\pi^{2}h^{2}}{(\pi^{2}-4C_{\mu}^{2})^{2}}.$
Hence, we obtain
$|e_{2}^{h}|_{1}^{2}\leq\frac{\pi^{2}h^{2}}{(\pi^{2}-4C_{\mu}^{2})^{2}}\cdot\frac{1}{2}\sum_{(\xi/\pi)=N}^{\infty}|\hat{f}(\xi)|^{2}.$
We define the quantities (and similarly $u_{\mathrm{low}}$ and $u_{\mathrm{high}}=e_{2}^{h}$)
$f_{\mathrm{low}}:=\sum_{(\xi/\pi)=1}^{N-1}\hat{f}(\xi)\sin(\xi x),\quad f_{\mathrm{high}}:=\sum_{(\xi/\pi)=N}^{\infty}\hat{f}(\xi)\sin(\xi x),\quad\|f\|:=\int_{0}^{1}|f(x)|^{2}~\mathrm{d}x.$ (6)
Then, it follows that
$|e_{2}^{h}|_{1}\leq\frac{\pi h}{\pi^{2}-4C_{\mu}^{2}}\|f_{\mathrm{high}}\|.$
Using higher regularity $f\in H_{0}^{p}(0,1)$ allows the split $\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}\frac{1}{\xi^{2p}}|\xi^{p}\hat{f}(\xi)|^{2}$ in (5) and yields
$|e_{2}^{h}|_{1}\leq\frac{h^{p+1}}{(\pi^{2}-4C_{\mu}^{2})\pi^{p-1}}|f_{\mathrm{high}}|_{p},\quad|f|_{p}:=\|f^{(p)}\|\text{ with $f^{(p)}$ the $p$th derivative},$ (7)
and also the $L^{2}$-norm of the downsampling error estimate
$\|e_{2}^{h}\|\leq\frac{h^{p+2}}{(\pi^{2}-4C_{\mu}^{2})\pi^{p}}|f_{\mathrm{high}}|_{p}.$ (8)
Remark 1.
Note that the separation $f=f_{\mathrm{low}}+f_{\mathrm{high}}$ depends on the mesh size $h=1/N$. We have $f_{\mathrm{low}}\to f$ and $f_{\mathrm{high}}\to 0$ in $H^{p}(0,1)$ as $N\to\infty$. By Parseval’s identity or orthogonality of the basis, we see that $|f|_{l}^{2}=|f_{\mathrm{low}}|_{l}^{2}+|f_{\mathrm{high}}|_{l}^{2}$ for all $l=0,..,p$. If $|f_{\mathrm{high}}|_{p}=O(k^{p})$ and $kh=O(1)$, then (7) and (8) say $|e_{2}^{h}|_{1}=O(h)$ and $\|e_{2}^{h}\|=O(h^{2})$, so the higher regularity does not improve the estimates.
The estimation of $e^{h}_{1}$ in case (i) and the error in case (ii) both depend on the particular finite difference schemes, which we shall discuss as follows.
3 Analysis of classical 3-point centred scheme
The scheme uses the approximation $u^{\prime\prime}(x)\approx(u(x+h)-2u(x)+u(x-h))/h^{2}$ to find $u^{h}$ satisfying
$(k^{2}-\frac{2}{h^{2}})u^{h}(x_{j})+\frac{1}{h^{2}}u^{h}(x_{j-1})+\frac{1}{h^{2}}u^{h}(x_{j+1})=f(x_{j}),\quad u^{h}(x_{0})=g_{0},\quad u^{h}(x_{N})=g_{1}.$ (9)
In case (i), $g_{0}=g_{1}=0$, applying the discrete sine transform to the above equation gives
$\left(k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}\right)\widehat{u^{h}}(\xi)=\widehat{f^{h}}(\xi)=\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi),\;\;\xi/\pi=1,..,N-1.$
In case (ii), $f=0$, the solution of (9) can be found in closed form,
$u^{h}(x_{j})=\frac{g_{1}-g_{0}\cos k^{h}}{\sin k^{h}}\sin(k^{h}x_{j})+g_{0}\cos(k^{h}x_{j}),\quad x_{j}=jh,\;j=0,1,..,N,$ (10)
where $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ is the discrete wavenumber. The solution is valid if and only if $k^{h}\not\in\mathbb{Z}\pi$ which we assume to hold. The formula (10) allows the grid function $u^{h}$ to be densely defined in $[0,1]$. Then the error $e^{h}(x):=u(x)-u^{h}(x)$ can be explicitly written as
$e^{h}(x)\!=\!A_{0}(x)g_{0}\!+\!A_{1}(x)g_{1},\quad A_{1}(x)\!=\!\frac{\sin(kx)}{\sin k}\!-\!\frac{\sin(k^{h}x)}{\sin k^{h}},\quad A_{0}(x)=A_{1}(1-x).$ (11)
We shall first analyze the case (ii), $f=0$, in subsection 3.1, and then case (i), $g_{0}=g_{1}=0$, in the remaining subsections.
3.1 Error for the Dirichlet problem with zero source
From (11), we have
$\|e^{h}\|\leq\|A_{0}\|\cdot|g_{0}|+\|A_{1}\|\cdot|g_{1}|=\|A_{1}\|(|g_{0}|+|g_{1}|).$
This inequality is sharp by considering $g_{0}=0$ or $g_{1}=0$. So the key is to estimate $\|A_{1}\|$. We calculate
$\int_{0}^{1}\frac{\sin^{2}(kx)}{\sin^{2}k}~\mathrm{d}x=\frac{1}{\sin^{2}k}\int_{0}^{1}\frac{1-\cos(2kx)}{2}~\mathrm{d}x=\frac{1}{2\sin^{2}k}\left(1-\frac{1}{2k}\sin(2k)\right),$
$\int_{0}^{1}\frac{\sin^{2}(k^{h}x)}{\sin^{2}k^{h}}~\mathrm{d}x=\frac{1}{2\sin^{2}k^{h}}\left(1-\frac{1}{2k^{h}}\sin(2k^{h})\right),$
$\int_{0}^{1}\frac{\sin(kx)\sin(k^{h}x)}{(\sin k)\sin k^{h}}~\mathrm{d}x$ $=\frac{1}{(\sin k)\sin k^{h}}\frac{1}{2}\int_{0}^{1}\cos((k-k^{h})x)-\cos((k+k^{h})x)~\mathrm{d}x$
$=\frac{1}{2(\sin k)\sin k^{h}}\left(\frac{\sin(k^{h}-k)}{k^{h}-k}-\frac{\sin(k+k^{h})}{k+k^{h}}\right).$
Therefore, we have
$\|A_{1}\|^{2}=$ $\int_{0}^{1}\left[\frac{\sin(kx)}{\sin k}-\frac{\sin(k^{h}x)}{\sin k^{h}}\right]^{2}~\mathrm{d}x$
$=$ $\frac{1}{2\sin^{2}k}+\frac{1}{2\sin^{2}k^{h}}-\frac{\cos k}{2k\sin k}-\frac{\cos k^{h}}{2k^{h}\sin k^{h}}-\frac{1}{(\sin k)\sin k^{h}}\left(\frac{\sin(k^{h}-k)}{k^{h}-k}-\frac{\sin(k+k^{h})}{k+k^{h}}\right)$
$=$ $\frac{1}{(\sin k)\sin k^{h}}\left[\frac{1}{2}\frac{\sin k^{h}}{\sin k}\!+\!\frac{1}{2}\frac{\sin k}{\sin k^{h}}\!-\!\frac{(\sin k^{h})\cos k}{2k}\!-\!\frac{(\sin k)\cos k^{h}}{2k^{h}}\!-\!\frac{\sin(k^{h}\!-\!k)}{k^{h}\!-\!k}\!+\!\frac{\sin(k\!+\!k^{h})}{k\!+\!k^{h}}\right]$
$=:$ $\frac{1}{(\sin k)\sin k^{h}}(S_{1}\!+\!S_{2}),$
where we separated the terms in the brackets into two parts
$S_{1}:=\frac{1}{2}\frac{\sin k^{h}}{\sin k}+\frac{1}{2}\frac{\sin k}{\sin k^{h}}-\frac{\sin(k^{h}-k)}{k^{h}-k},\quad S_{2}:=-\frac{(\sin k^{h})\cos k}{2k}-\frac{(\sin k)\cos k^{h}}{2k^{h}}+\frac{\sin(k+k^{h})}{k+k^{h}},$ (12)
where $S_{1}$ mainly accounts for the dispersion error, and $S_{2}$ is small for large $k$. Note that
$\frac{\sin k^{h}}{\sin k}$ $=\frac{\sin(k^{h}-k+k)}{\sin k}=\frac{\sin(k^{h}-k)\cos k}{\sin k}+\cos(k^{h}-k),$
$\frac{\sin k}{\sin k^{h}}$ $=\frac{\sin(k-k^{h}+k^{h})}{\sin k^{h}}=\frac{\sin(k-k^{h})\cos k^{h}}{\sin k^{h}}+\cos(k-k^{h}).$
So we have
$S_{1}=\frac{1}{2}\sin(k^{h}-k)\left(\cot k-\cot k^{h}\right)+\cos(k^{h}-k)-\frac{\sin(k^{h}-k)}{k^{h}-k}.$ (13)
We have also
$S_{2}=$ $-\frac{(\sin k^{h})\cos k}{2k}-\frac{(\sin k)\cos k^{h}}{2k^{h}}+\frac{(\sin k)\cos k^{h}+(\sin k^{h})\cos k}{k+k^{h}}$
$=$ $\frac{k^{h}-k}{2(k+k^{h})}\left[\frac{(\sin k)\cos k^{h}}{k^{h}}-\frac{(\sin k^{h})\cos k}{k}\right]$
$=$ $\frac{k^{h}-k}{2(k+k^{h})}\left[\frac{(\sin k)\cos k^{h}-(\sin k^{h})\cos k}{k}+\left(\frac{1}{k^{h}}-\frac{1}{k}\right)(\sin k)\cos k^{h}\right]$
$=$ $\frac{k^{h}-k}{2(k+k^{h})}\left[\frac{\sin(k-k^{h})}{k}+\frac{k-k^{h}}{kk^{h}}(\sin k)\cos k^{h}\right].$ (14)
Lemma 1.
Let $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$. Then $|\sin k|\geq\frac{2}{\pi}\sigma_{k}.$
Proof.
Let $n_{k}\in\mathbb{N}$ be such that $|k-n_{k}\pi|\leq\frac{\pi}{2}$. Since $\sin x\geq\frac{2}{\pi}x$ for $0\leq x\leq\frac{\pi}{2}$, we have
$|\sin k|=\sin|k-n_{k}\pi|\geq\frac{2}{\pi}|k-n_{k}\pi|\geq\frac{2}{\pi}\sigma_{k},$
where in the last step we used the definition of $\sigma_{k}$. ∎
Lemma 2.
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $c>0$. If $0<\frac{kh}{2}<1$, then $\frac{1}{24}k^{3}h^{2}<k^{h}-k<\frac{\pi-2}{8}k^{3}h^{2}$. If in addition $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$ and $k^{3}h^{2}<\frac{8c\sigma_{k}}{\pi-2}$, then $k^{h}-k<c\sigma_{k}$ and $|\sin k^{h}|>\frac{2(1-c)}{\pi}\sigma_{k}$.
Proof.
Since $x+\frac{x^{3}}{6}<\arcsin x<x+(\frac{\pi}{2}-1)x^{3}$ for $0<x<1$, we have $k\!+\!\frac{1}{24}k^{3}h^{2}<k^{h}<k\!+\!\frac{\pi\!-\!2}{8}k^{3}h^{2}<k\!+\!c\sigma_{k}$. It follows that $\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k^{h}|>\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k|-|k^{h}\!-\!k|>(1\!-\!c)\sigma_{k}$. So $|\sin k^{h}|\geq\frac{2(1\!-\!c)}{\pi}\sigma_{k}$. ∎
Remark 2.
It is important to require $k^{3}h^{2}$ being small for convergence of the finite difference scheme. Otherwise, Lemma 2 says that $k^{h}-k$ can be large. For example, when $k$ is large and $k^{h}-k=\pi$, we have $S_{1}=-1$, $S_{2}<\frac{\pi^{2}}{4k^{3}}$ and $\|A_{1}\|>1-\frac{\pi^{2}}{4k^{3}}$.
Lemma 3.
Under the assumptions of Lemma 2 with $0<c<1$, for $S_{1}$ in (13) it holds that
$\frac{1}{8}(k^{h}-k)^{2}<S_{1}<\left(\frac{\pi^{2}}{4\sigma_{k}^{2}}+\frac{1}{6}\right)(k^{h}-k)^{2}.$ (15)
Proof.
Let $k\in((n_{k}^{+}-1)\pi,n_{k}^{+}\pi)$ for an integer $n_{k}^{+}$. By Lemma 2, it holds that $0<k^{h}-k<c\sigma_{k}<\frac{\pi}{2}$. Moreover, $(n_{k}^{+}-1)\pi+\sigma_{k}\leq k<k^{h}<k+c\sigma_{k}\leq n_{k}^{+}\pi-(1-c)\sigma_{k}$. Hence $(n_{k}^{+}-1)\pi<k<k^{h}<n_{k}^{+}\pi$, and $\cot k-\cot k^{h}=\frac{k^{h}-k}{\sin^{2}\tilde{k}}>k^{h}-k$ for some $\tilde{k}\in(k,k^{h})$. Since
[…]
it can be deduced from (13) that
$S_{1}<\frac{1}{2}\frac{(k^{h}-k)^{2}}{\sin^{2}k}+1-\left(1-\frac{1}{6}(k^{h}-k)^{2}\right).$
After simplification, the upper bound in (15) is obtained. The lower bound is derived from (12),
$S_{1}>1-\left(1-\frac{1}{8}(k^{h}-k)^{2}\right)=\frac{1}{8}(k^{h}-k)^{2},$
where $0<k^{h}-k<\frac{\pi}{2}$, $(n_{k}^{+}-1)\pi<k<k^{h}<n_{k}^{+}\pi$ and $x+\frac{1}{x}\geq 2$ for $x>0$ are used. ∎
Lemma 4.
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $S_{2}$ given in (14). If $0<\frac{kh}{2}<1$, then it holds that
$\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<S_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$
Proof.
The conclusion follows directly from $k^{h}>k>0$, $|\sin x|\leq|x|$, $|\sin x|\leq 1$ and $|\cos x|\leq 1$. ∎
Theorem 1 ($L^{2}$ Error for $f=0$).
Under the assumptions of Lemma 2, for the $L^{2}$-norm of $A_{1}$ in (11), it holds that
$\frac{1}{24}\sqrt{\frac{1}{8}-\frac{1}{4k^{2}}-\frac{1}{4k^{3}}}k^{3}h^{2}<\|A_{1}\|<\frac{\pi(\pi-2)}{16\sigma_{k}\sqrt{1-c}}\sqrt{\frac{\pi^{2}}{4\sigma_{k}^{2}}+\frac{1}{6}+\frac{1}{4k^{2}}+\frac{1}{4k^{3}}}k^{3}h^{2},$
where $\sigma_{k}=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$ and $k^{3}h^{2}\frac{(\pi-2)}{8\sigma_{k}}<c<1$. Moreover, the error $e^{h}=u-u^{h}$ satisfies $\|e^{h}\|\leq\|A_{1}\|(|g_{0}|+|g_{1}|)$ with equality attained when $g_{0}=0$ or $g_{1}=0$, and the relative error satisfies
$\frac{\|e^{h}\|}{\|u\|}\leq\frac{\sqrt{2}|\sin k|}{\sqrt{1-\frac{\sin 2k}{2k}}}\|A_{1}\|$
with equality attained when $g_{0}=0$ or $g_{1}=0$.
Proof.
The absolute error estimate is obtained by combining Lemmas 1, 2, 3 and 4. The relative error estimate follows from
$|u\|\geq\left||g_{1}|\cdot\left\|\frac{\sin(kx)}{\sin k}\right\|-|g_{0}|\cdot\left\|\frac{\sin(k(1-x))}{\sin k}\right\|\right$|
with equality attained when $g_{0}=0$ or $g_{1}=0$. ∎
It is also interesting to estimate the $H^{1}$-semi-norm of $A_{1}(x)$ in (11). Some calculations lead to
$|A_{1}|_{1}^{2}=\int_{0}^{1}\left(k\frac{\cos(kx)}{\sin k}-k^{h}\frac{\cos(k^{h}x)}{\sin k^{h}}\right)^{2}~\mathrm{d}x=\frac{kk^{h}}{\sin k\sin k^{h}}(\tilde{S}_{1}+\tilde{S}_{2}),$
where we introduced
$\tilde{S}_{1}:=\frac{1}{2}\frac{k\sin k^{h}}{k^{h}\sin k}+\frac{1}{2}\frac{k^{h}\sin k}{k\sin k^{h}}-\frac{\sin(k^{h}-k)}{k^{h}-k},\quad\tilde{S}_{2}:=\frac{(\sin k^{h})\cos k}{2k^{h}}+\frac{(\sin k)\cos k^{h}}{2k}-\frac{\sin(k+k^{h})}{k+k^{h}}.$ (16)
More calculations as previously done for $S_{1}$ and $S_{2}$ yield
$\tilde{S}_{1}=\frac{1}{2}\sin(k^{h}-k)\left(\cot k-\cot k^{h}\right)+\cos(k^{h}-k)-\frac{\sin(k^{h}-k)}{k^{h}-k}\\ \qquad+(k^{h}-k)\left[-(\sin(k^{h}-k))\left(\frac{\cot k}{2k^{h}}+\frac{\cot k^{h}}{2k}\right)+(\cos(k^{h}-k))\frac{k^{h}-k}{2kk^{h}}\right],$ (17)
$\tilde{S}_{2}=\frac{k^{h}-k}{2(k+k^{h})}\left[\frac{\sin(k-k^{h})}{k}+\frac{k^{h}-k}{kk^{h}}(\sin k^{h})\cos k\right].$ (18)
Lemma 5.
Under the assumptions of Lemma 2 with $0<c<1$, for $\tilde{S}_{1}$ in (17) it holds that
$\frac{1}{8}(k^{h}-k)^{2}<\tilde{S}_{1}<\left(\frac{\pi^{2}}{4\sigma_{k}^{2}}+\frac{1}{6}+\frac{\pi(2-c)}{4\sigma_{k}(1-c)k}+\frac{1}{2k^{2}}\right)(k^{h}-k)^{2}.$
Proof.
The lower bound is derived from (16) similarly to the proof of Lemma 3. For the upper bound, note that the first line of (17) equals $S_{1}$, and the remaining line satisfies
$\,(k^{h}-k)\left[-(\sin(k^{h}-k))\left(\frac{\cot k}{2k^{h}}+\frac{\cot k^{h}}{2k}\right)+(\cos(k^{h}-k))\frac{k^{h}-k}{2kk^{h}}\right]$
$<$ $\,(k^{h}-k)^{2}\left(\frac{\pi(2-c)}{4\sigma_{k}(1-c)k}+\frac{1}{2k^{2}}\right),$
which, combined with the upper bound for $S_{1}$ in Lemma 3, gives the conclusion. ∎
Lemma 6.
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $\tilde{S}_{2}$ be given in (18). If $0<\frac{kh}{2}<1$, then it holds that
$\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<\tilde{S}_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$
Proof.
The conclusion follows directly from $k^{h}>k>0$, $|\sin x|\leq|x|$, $|\sin x|\leq 1$ and $|\cos x|\leq 1$. ∎
Theorem 2 ($H^{1}$ error for $f=0$).
Under the assumptions of Lemma 2, for the $H^{1}$-semi-norm of $A_{1}$ in (11), it holds that
$\frac{1}{24}\sqrt{\frac{1}{8}-\frac{1}{4k^{2}}-\frac{1}{4k^{3}}}k^{4}h^{2}<|A_{1}|_{1}<\frac{\pi(\pi-2)}{16\sigma_{k}\sqrt{1-c}}\sqrt{\frac{\pi^{2}}{2\sigma_{k}^{2}}+\frac{1}{3}+\frac{\pi(2-c)}{2\sigma_{k}(1-c)k}+\frac{3}{2k^{2}}+\frac{1}{2k^{3}}}k^{4}h^{2},$
where $\sigma_{k}=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$ and $k^{3}h^{2}\frac{(\pi-2)}{8\sigma_{k}}<c<1$. Moreover, the error $e^{h}=u-u^{h}$ satisfies $|e^{h}|_{1}\leq|A_{1}|_{1}(|g_{0}|+|g_{1}|)$ with equality attained when $g_{0}=0$ or $g_{1}=0$, and the relative error satisfies
$\frac{|e^{h}|_{1}}{|u|_{1}}\leq\frac{\sqrt{2}|\sin k|}{\sqrt{1+\frac{\sin 2k}{2k}}}\cdot\frac{|A_{1}|_{1}}{k}$
with equality attained when $g_{0}=0$ or $g_{1}=0$.
Proof.
The conclusion is obtained by combining Lemmas 1, 2, 5, 6, and $k^{h}=\frac{2}{h}\arcsin\frac{kh}{2}<\frac{\pi}{2}k<2k$. ∎
3.2 Well-posedness of the classical 3-point centred scheme
The discretized problem is well-posed if and only if $\lambda^{h}(\xi):=k^{2}\!\!-\!\!\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}\neq 0$ for $\xi\in\{1,..,N\!\!-\!\!1\}\pi$. Given that the continuous problem (1) is well-posed, we can expect for all sufficiently small $h$ that the discretized problem is also well-posed. A corresponding upper bound for $h$ is:
Lemma 7.
Suppose $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$. Let $n_{k}^{+}\in\mathbb{N}$ such that $(n_{k}^{+}\!\!-\!\!1)\pi<k<n_{k}^{+}\pi$. If
$0<h=\frac{1}{N}<h_{k}:=2\sqrt{3\sigma_{k}}/\left(\frac{\sigma_{k}}{2}+n_{k}^{+}\pi\right)^{3/2}=O(\sqrt{\sigma_{k}}k^{-3/2}),$ (19)
then
$\sigma_{k}^{h}:=\min_{(\xi/\pi)=1,..,N-1}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|\geq\frac{\sigma_{k}}{2},$ (20)
and $\lambda^{h}:=k^{2}\!\!-\!\!\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}\geq\max\{2k-\frac{\sigma_{k}}{2},k+\frac{2}{\pi}\}\frac{\sigma_{k}}{2}$.
Proof.
If $\frac{2}{h}\sin\frac{\xi h}{2}\geq n_{k}^{+}\pi$ or $\frac{2}{h}\sin\frac{\xi h}{2}\leq(n_{k}^{+}-1)\pi$, then $\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|\geq\sigma_{k}$. If $\xi<k$, then
$\frac{2}{h}\sin\frac{\xi h}{2}\leq\xi<k\text{ and }\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|=k-\frac{2}{h}\sin\frac{\xi h}{2}\geq k-\xi=|k-\xi|\geq\sigma_{k}.$
We need only to consider the case $(n_{k}^{+}-1)\pi<\frac{2}{h}\sin\frac{\xi h}{2}<n_{k}^{+}\pi$ and $\xi\geq n_{k}^{+}\pi$. It suffices to show that $n_{k}^{+}\pi-\frac{2}{h}\sin\frac{\xi h}{2}<\frac{\sigma_{k}}{2}$, which is equivalent to
$\frac{\xi h}{2}-\sin\frac{\xi h}{2}<\frac{\sigma_{k}h}{4}+\frac{\xi h-n_{k}^{+}\pi h}{2}.$
It is thus enough to show that $\theta-\sin\theta<\sigma_{k}h/4$ for $\theta=\xi h/2$. Since $\sin\theta>\theta-\frac{\theta^{3}}{3!}$ for $\theta\in(0,\pi/2)$, it suffices to show $\frac{\theta^{3}}{3!}<\sigma_{k}h/4$, which is equivalent to $\sin\theta<\sin\sqrt[3]{{3}\sigma_{k}h/2}=:\sin v$. Since $\sin\theta<n_{k}^{+}\pi h/2$ and $\sin v>v-\frac{v^{3}}{3!}$, it is enough to show that $n_{k}^{+}\pi h/2<v-\frac{v^{3}}{3!}$, which is
${n_{k}^{+}\pi h}/{2}<\sqrt[3]{3\sigma_{k}h/2}-(\sigma_{k}h/4).$
Solving the inequality for $h$ gives the conclusion. ∎
Remark 3.
If we define $h_{k}^{*}>0$ as the largest reciprocal of an integer such that
$\min_{(\xi/\pi)=1,..,N-1}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|\geq\frac{\sigma_{k}}{2}$ (21)
for all $h=1/N\in(0,h_{k}^{*})$, then $h_{k}^{*}$ is likely of the same order in $k$ as the upper bound $h_{k}$ in (19) is. This is illustrated on the left of Fig. 1. But Lemma 7 does not prohibit using larger mesh size. The existence of a mesh size $h\sim k^{-1}$ that verifies $\sigma_{k}^{h}\geq\sigma_{k}/2$ is shown on the right of Fig. 1.
Figure 1: Admissible mesh size $h$ for well-posedness of (9). Left: Upper bounds $h_{k}$ (19) and $h_{k}^{*}$ (21) so that all smaller $h$ verifies (20) with $\sigma_{k}=1$. Right: Existence of $h=O(k^{-1})$ verifying (20) with $\sigma_{k}=1$.
3.3 Aliasing error of the classical 3-point centred scheme
From now on, we assume that $\lambda^{h}\neq 0$. Then the derivative of the partial error $e_{1}^{h}$ in (3) is
$(e_{1}^{h})^{\prime}(x)$ $=\sum_{(\xi/\pi)=1}^{N-1}\left(\hat{u}(\xi)-\widehat{u^{h}}(\xi)\right)\xi\cos(\xi x)=(E_{1}^{h})^{\prime}(x)-(E_{2}^{h})^{\prime}(x)$ (22)
$=\sum_{(\xi/\pi)=1}^{N-1}\left(\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right)\hat{f}(\xi)\cos(\xi x)-\sum_{(\xi/\pi)=1}^{N-1}\frac{\xi}{\lambda^{h}}\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\cos(\xi x).$
To estimate $E_{2}^{h}$, we first note that
$\left|\sum_{m=1}^{\infty}\hat{f}(\pm\xi+2mN\pi)\right|\leq\sqrt{\sum_{m=1}^{\infty}\frac{1}{(\pm\xi+2mN\pi)^{2p}}}\cdot\sqrt{\sum_{m=1}^{\infty}\left|(\pm\xi+2mN\pi)^{p}\hat{f}(\pm\xi+2mN\pi)\right|^{2}},$ (23)
and for $0\leq\xi\leq(N-1)\pi$ that
$\sum_{m=1}^{\infty}\frac{1}{(\pm\xi+2mN\pi)^{2p}}\leq\frac{1}{(\pm\xi+2N\pi)^{2p}}+\frac{1}{2N\pi}\int_{\pm\xi+2N\pi}^{\infty}\frac{1}{t^{2p}}~\mathrm{d}t\leq\frac{2p}{2p-1}\frac{1}{((\frac{3}{2}\pm\frac{1}{2})N\pi)^{2p}}.$ (24)
We can then estimate
$\;|E_{2}^{h}|_{1}^{2}=\sum_{(\xi/\pi)=1}^{N-1}\frac{\xi^{2}}{\left|\lambda^{h}\right|^{2}}\left|\sum_{s=\pm 1}\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\right|^{2}$
$\leq$ $\;\max_{\xi\in\{1,..N-1\}\pi}\frac{\xi^{2}}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left|\sum_{s=\pm 1}\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\right|^{2}$
$\leq$  $2\max_{\xi\in\{1,..N-1\}\pi}\frac{\xi^{2}}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left[\left|\sum_{m=1}^{\infty}\hat{f}(\xi+2mN\pi)\right|^{2}+\left|\sum_{m=1}^{\infty}\hat{f}(-\xi+2mN\pi)\right|^{2}\right].$
Therefore, for $p\geq 1$, using (24), we have
$|E_{2}^{h}|_{1}\leq 2\max_{\xi\in\{1,..N-1\}\pi}\frac{\xi}{|\lambda^{h}|}\frac{h^{p}}{\pi^{p}}|f_{\mathrm{high}}|_{p}.$ (25)
Lemma 8.
Suppose $k>\pi$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
Then, for $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$, the $H^{1}$-semi-norm of the aliasing error $E_{2}^{h}$ in (22) satisfies
$|E_{2}^{h}|_{1}\leq\frac{1}{\tilde{\sigma}_{k}\pi^{p-1}}h^{p}|f_{\mathrm{high}}|_{p}.$ (26)
Proof.
Based on (25), it suffices to find an upper bound of
$\frac{\xi}{|\lambda^{h}|}=\frac{\xi}{|k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}|}\quad\text{ over }\xi\in\{1,\ldots,N-1\}\pi.$
Denote $\theta_{k}:=\arcsin\frac{kh}{2}$ and $\theta:=\frac{\xi h}{2}$. Then
$\frac{\xi}{|\lambda^{h}|}=\frac{h}{2}\frac{\theta}{|\sin^{2}\theta_{k}-\sin^{2}\theta|}=:\frac{h}{2}\phi(\theta)\quad\text{for }\theta\in\{1,\ldots,N-1\}\frac{\pi h}{2}=:\Theta.$
When $\theta<\theta_{k}$, the function $\phi(\theta)$ increases with $\theta$. Let $\theta_{-}:=\max\{\theta\in\Theta:~\theta<\theta_{k}\}$. So
$\phi(\theta)\leq\frac{\theta_{-}}{\sin^{2}\theta_{k}-\sin^{2}\theta_{-}}\leq\frac{2}{h\tilde{\sigma}_{k}}\cdot\frac{\theta_{-}}{\sin\theta_{k}+\sin\theta_{-}}\leq\frac{2}{h\tilde{\sigma}_{k}}\cdot\frac{\theta_{-}}{\frac{2}{\pi}\theta_{k}+\frac{2}{\pi}\theta_{-}}<\frac{\pi}{2h\tilde{\sigma}_{k}}\quad\text{when }\theta\in\Theta\cap(0,\theta_{k}).$
When $\theta>\theta_{k}$, the function $\phi(\theta)$ has the derivative
$\phi^{\prime}(\theta)=\frac{\sin^{2}\theta-\sin^{2}\theta_{k}-\theta\sin(2\theta)}{(\sin^{2}\theta-\sin^{2}\theta_{k})^{2}}.$
So any stationary point $\theta_{c}$ of $\phi$ satisfies $\sin^{2}\theta_{c}-\sin^{2}\theta_{k}-\theta_{c}\sin(2\theta_{c})=0$ and
$\phi(\theta_{c})=\frac{\theta_{c}}{\sin^{2}\theta_{c}-\sin^{2}\theta_{k}}=\frac{1}{\sin(2\theta_{c})}.$
Le $\theta_{+}:=\min\{\theta\in\Theta:~\theta>\theta_{k}\}$ and $\theta_{\max}:=\max\Theta$. If $\theta_{+}<\frac{\pi}{4}$, then the numerator of $\phi^{\prime}(\theta_{+})$ satisfies
$\sin^{2}\theta_{+}-\sin^{2}\theta_{k}-\theta_{+}\sin(2\theta_{+})=\sin(2\tilde{\theta})(\theta_{+}-\theta_{k})-\theta_{+}\sin(2\theta_{+})<0,$
because by the mean value theorem $\tilde{\theta}<\theta_{+}$. Note that the derivative of the numerator of $\phi^{\prime}(\theta)$ is $2\theta\cos(2\theta)$. So if $\theta_{+}<\frac{\pi}{4}$, then $\phi^{\prime}<0$ on $\theta_{+},\frac{\pi}{4}]$. Therefore, no matter $\theta_{+}<\frac{\pi}{4}$ or not, any stationary point $\theta_{c}$ of $\phi$ in $\theta_{+},\theta_{\max}]$ must satisfy $\theta_{c}\geq\frac{\pi}{4}$. It follows that
$\phi(\theta_{c})\leq\frac{1}{\sin(2\theta_{\max})}=\frac{1}{\sin((N-1)\pi h)}=\frac{1}{\sin(\pi h)}\leq\frac{1}{2h}.$
Since $k>\pi$, the numerator of $\phi^{\prime}(\theta_{\max})$ satisfies
$\sin^{2}\frac{\pi h}{2}-\sin^{2}\theta_{k}-(1-h)\frac{\pi}{2}\sin(\pi h)<\frac{\pi^{2}h^{2}}{4}-\frac{k^{2}h^{2}}{4}-(1-h)\pi h<0.$
Note that
$\phi(\theta_{+})=\frac{\theta_{+}}{\sin^{2}\theta_{k}-\sin^{2}\theta_{+}}\leq\frac{2}{h\tilde{\sigma}_{k}}\cdot\frac{\theta_{+}}{\sin\theta_{k}+\sin\theta_{+}}<\frac{2}{h\tilde{\sigma}_{k}}\cdot\frac{\theta_{+}}{\sin\theta_{+}}<\frac{\pi}{h\tilde{\sigma}_{k}}.$
So we obtain
$\max_{\theta\in\Theta}\phi(\theta)<\max\{\frac{\pi}{h\tilde{\sigma}_{k}},\frac{1}{2h}\}=\frac{\pi}{h\tilde{\sigma}_{k}},\quad\text{and }\max_{\xi\in\{1,\ldots,N-1\}}\frac{\xi}{|\lambda^{h}|}<\frac{\pi}{2\tilde{\sigma}_{k}}.$
Substituting this into (25) gives the conclusion. ∎
Remark 4.
By Lemma 8 the second part $(E_{2}^{h})^{\prime}$ of $(e_{1}^{h})^{\prime}$ in (22) is controlled, similarly to $(e_{2}^{h})^{\prime}$ in (4), by the grid under-resolved source $f_{\mathrm{high}}$, and behaves as $O(h^{p})$ when $k>\pi$, $\frac{kh}{2}\leq C_{\mu}<1$ and $\tilde{\sigma}_{k}>0$.
The $L^{2}$-norm of the aliasing error can also be derived. Recall that
$e_{1}^{h}(x)$ $=\sum_{(\xi/\pi)=1}^{N-1}\left(\hat{u}(\xi)-\widehat{u^{h}}(\xi)\right)\sin(\xi x)=E_{1}^{h}(x)-E_{2}^{h}(x)$ (27)
$=\sum_{(\xi/\pi)=1}^{N-1}\left(\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right)\hat{f}(\xi)\sin(\xi x)-\sum_{(\xi/\pi)=1}^{N-1}\frac{1}{\lambda^{h}}\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\sin(\xi x).$
Lemma 9.
Suppose $k>\pi$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
Then, for $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$, the $L^{2}$-norm of the aliasing error $E_{2}^{h}$ in (27) satisfies
$\|E_{2}^{h}\|\leq\frac{2}{\tilde{\sigma}_{k}k\pi^{p}}h^{p}|f_{\mathrm{high}}|_{p}.$
Proof.
Based on (23) and (24), we have
$\;\|E_{2}^{h}\|^{2}=\sum_{(\xi/\pi)=1}^{N-1}\frac{1}{\left|\lambda^{h}\right|^{2}}\left|\sum_{s=\pm 1}\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\right|^{2}$
$\leq$ $\;\max_{\xi\in\{1,..N-1\}\pi}\frac{1}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left|\sum_{s=\pm 1}\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi)\right|^{2}$
$\leq$  $2\max_{\xi\in\{1,..N-1\}\pi}\frac{1}{|\lambda^{h}|^{2}}\sum_{(\xi/\pi)=1}^{N-1}\left[\left|\sum_{m=1}^{\infty}\hat{f}(\xi+2mN\pi)\right|^{2}+\left|\sum_{m=1}^{\infty}\hat{f}(-\xi+2mN\pi)\right|^{2}\right].$
Therefore, for $p\geq 1$, using (24), we get
$\|E_{2}^{h}\|\leq 2\max_{\xi\in\{1,\ldots,N-1\}\pi}\frac{1}{|\lambda^{h}|}\frac{h^{p}}{\pi^{p}}|f_{\mathrm{high}}|_{p}.$
Note that
$\frac{1}{|\lambda^{h}|}=\frac{1}{|k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}|}=\frac{1}{|k-\frac{2}{h}\sin\frac{\xi h}{2}|}\frac{1}{|k+\frac{2}{h}\sin\frac{\xi h}{2}|}\leq\frac{1}{\tilde{\sigma}_{k}k}.$
Combining the above two inequalities gives the conclusion. ∎
Remark 5.
By Lemma 9, the $L^{2}$-norm of the aliasing error is $O\left(\frac{h^{p}}{\tilde{\sigma}_{k}k}\right)$. This can be compared to the $H^{1}$-semi-norm of the aliasing error which is $O\left(\frac{h^{p}}{\tilde{\sigma}_{k}}\right)$; see Lemma 8.
3.4 Absolute error estimates of the classical 3-point centred scheme
We now turn to estimate the first part of $(e_{1}^{h})^{\prime}$ given in (3) and (22),
$(E_{1}^{h})^{\prime}(x)=\sum_{(\xi/\pi)=1}^{N-1}\left(\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right)\hat{f}(\xi)\cos(\xi x),$ (28)
the only part of $(e^{h})^{\prime}=(E_{1}^{h})^{\prime}-(E_{2}^{h})^{\prime}+(e_{2}^{h})^{\prime}$ controlled by the grid well-resolved source $f_{\mathrm{low}}$. Since
$|E_{1}^{h}|_{1}^{2}=\frac{1}{2}\sum_{(\xi/\pi)=1}^{N-1}\left|\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right|^{2}|\hat{f}(\xi)|^{2}\leq\|f_{\mathrm{low}}\|^{2}\max_{\xi\in\{1,..,N-1\}\pi}\left|\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right|^{2},$ (29)
we need only to find the maximum of
$\psi(\xi):=\left|\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right|=\left|\frac{\xi}{k^{2}-\xi^{2}}-\frac{\xi}{k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}}\right|=\frac{h}{2}\left|\frac{\theta}{\mu^{2}-\theta^{2}}-\frac{\theta}{\mu^{2}-\sin^{2}\theta}\right|=:\frac{h}{2}\phi(\theta)$ (30)
for $\mu:=kh/2$ and $\theta:=\xi h/2$.
Remark 6.
The inequality in (29) is sharp. Let $\xi_{*}:=\arg\max_{\xi\in\{1,..N-1\}\pi}\psi(\xi)$. Then the inequality holds with equality if $f_{\mathrm{low}}=\hat{f}(\xi_{*})\sin(\xi_{*}x)$. In the convergence test, one usually takes a fixed $f$ as $h\to 0$. So the sharpness may be perturbed if $\xi_{*}$ depends on $h$. This is not an issue, if $\xi_{*}$ is independent of $h$ or if $\xi_{*}$ varies slightly around a fixed number, which is the case for $\xi_{*}$ being the continuous frequencies $\xi$ closest to $k$, or their discrete counterparts $\frac{2}{h}\sin\frac{\xi h}{2}$ closest to $k$, as we shall see in Lemma 10. However, for the case $\xi_{*}=(N-1)\pi$ which we shall also see in Lemma 10, the sharpness will require a changing $f$ at higher and higher frequency as $h\to 0$. In that case, we will derive an alternate estimate to (29) using the smoothness of a fixed $f$.
Lemma 10.
Suppose $k>2\pi$, $\mu:=\frac{kh}{2}\leq C_{\mu}<1$, $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$, and there is a sequence of $h=\frac{1}{N}$ ($N\geq 4$) going to zero such that
[…]
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$, $\xi_{\max}=(N-1)\pi$, and $k_{\pm}\in\pi\mathbb{N}$ (resp. $k^{h}_{\pm}\in\pi\mathbb{N}$) be the closest points such that $k_{-}<k<k_{+}$ (resp. $k^{h}_{-}<k^{h}<k^{h}_{+}$). Then, $k_{+}<\xi_{\max}$, and
$\max_{\xi\in\{1,..,N-1\}\pi}\psi(\xi)=\max\{\psi(k_{-}),\psi(k_{+}),\psi(k_{-}^{h}),\psi(k_{+}^{h}),\psi(\xi_{\max})\},$
where $\psi$ is given in (30), $\psi(k_{\pm}^{h})$ can be removed if $k_{\pm}^{h}\geq\xi_{\max}$, and $k_{+}$ (resp. $k_{-}^{h}$) equals $k_{+}^{h}$ (resp. $k_{-}$) when $\pi\mathbb{N}\cap(k,k^{h})=\emptyset$. Moreover, we have
$\frac{k^{3}h^{2}}{(240\sigma_{-}+k^{3}h^{2})6\sigma_{-}}<$ $\,\psi(k_{-})<\frac{k^{3}h^{2}}{(24\sigma_{-}+k^{3}h^{2})2\sigma_{-}}\,$,
$\frac{1}{15}\frac{(k+\sigma_{+})^{3}h^{2}}{\sigma_{+}(4\pi+k^{3}h^{2})}<$ $\,\psi(k_{+})<\frac{2}{3}\frac{k^{3}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,$, $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{3}h^{2}}{12\pi(4\sigma_{+}+k^{3}h^{2})}<$ $\,\psi(k_{-}^{h})<\frac{\pi^{4}}{384}\frac{k^{3}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,$, $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{3}h^{2}}{12\pi(4\sigma_{+}^{h}+k^{3}h^{2})}<\,$ $\psi(k_{+}^{h})<\frac{(1+\pi)^{4}}{16}\frac{k^{3}h^{2}}{\sigma_{+}^{h}(24\sigma_{+}^{h}+k^{3}h^{2})\cos(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2})}\,$, $\text{ if }C_{\mu}<\cos\frac{\sigma_{+}^{h}h}{2},$
$\frac{h}{\pi}\left(\frac{9\pi^{2}}{64}\!-\!1\right)<$ $\,\psi(\xi_{\max})<\frac{4(\pi^{2}-4)\pi h}{\left(\cos^{2}\frac{\pi h}{2}\!-\!C_{\mu}^{2}\right)(9\pi^{2}-64)}\,$, $\text{ if }C_{\mu}<\cos\frac{\pi h}{2},$
where $\sigma_{\pm}:=|k-k_{\pm}|\geq\sigma_{k}=\min\{\sigma_{-},\sigma_{+}\}$, $\sigma_{\pm}^{h}:=|k^{h}-k_{\pm}^{h}|\geq\tilde{\sigma}_{k}$ and $\theta_{\mu}:=\arcsin C_{\mu}$.
Remark 7.
Note that $\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset$ if and only if $k_{+}<k^{h}$. Using $\mu^{3}/6<\arcsin\mu-\mu<(\pi-2)\mu^{3}/2$ for $0<\mu<1$, we find a sufficient condition for $k_{+}\geq k^{h}$ is that $k^{3}h^{2}\leq 8\sigma_{+}/(\pi-2)$ and a sufficient condition for $k_{+}<k^{h}$ is that $k^{3}h^{2}>24\sigma_{+}$.
Remark 8.
On the one hand, if $k^{3}h^{2}$ is greater than a positive constant, then the lower bound of $\psi(k_{-})$ is greater than a positive constant independent of $k$ and $h$, no matter how small $kh$ is. On the other hand, if $k^{3}h^{2}<C$, then based on Lemma 10 we can find the bounds
$\frac{k^{3}h^{2}}{(240\sigma_{-}+C)6\sigma_{-}}<\psi(k_{-})<\frac{k^{3}h^{2}}{48\sigma_{-}^{2}},$
which are exactly of order $k^{3}h^{2}$. For $\psi(k_{+})$, $\psi(k_{-}^{h})$ and $\psi(k_{+}^{h})$, we have similar results.
Proof of Lemma 10.
Let $\Theta:=\{1,..,N\!-\!1\}\frac{\pi h}{2}$, $\theta_{k}:=\arcsin\mu$, $\theta_{\max}:=\frac{h}{2}\xi_{\max}=\frac{\pi}{2}(1-h)$, $\mu_{\pm}:=\frac{h}{2}k_{\pm}$, and $\theta_{\pm}:=\frac{h}{2}k^{h}_{\pm}$. We first show that $\sigma_{\pm}^{h}\geq\tilde{\sigma}_{k}$. Indeed, by the mean value theorem, it holds
$|k^{h}-k_{\pm}^{h}|=\frac{2}{h}\left|\arcsin\frac{kh}{2}-\frac{k_{\pm}^{h}h}{2}\right|=\frac{2}{h}|\theta_{k}-\theta_{\pm}|\geq\frac{2}{h}\left|\sin\theta_{k}-\sin\theta_{\pm}\right|=\left|k-\frac{2}{h}\sin\frac{k_{\pm}^{h}h}{2}\right|.$
Recall the definition of $\phi$ in (30). When $0<\theta<\mu$, the function
$\phi(\theta)=\frac{(\theta^{2}-\sin^{2}\theta)\theta}{(\mu^{2}-\theta^{2})(\mu^{2}-\sin^{2}\theta)}$
is increasing because the numerator is increasing and the denominator is decreasing. Indeed, $(\theta^{2}-\sin^{2}\theta)^{\prime}=2\theta-\sin(2\theta)>0$. So the $\max\phi(\theta)$ over $0,\mu_{-}]$ is attained at $\theta=\mu_{-}$.
When $\mu<\theta<\theta_{k}$, we have
$\phi(\theta)=\frac{\theta}{\theta^{2}-\mu^{2}}+\frac{\theta}{\mu^{2}-\sin^{2}\theta}=\frac{1}{2}\left(\frac{1}{\theta-\mu}+\frac{1}{\theta+\mu}\right)+\frac{1}{2\mu}\left(\frac{\theta}{\mu-\sin\theta}+\frac{\theta}{\mu+\sin\theta}\right).$
Note that $1/(\theta\pm\mu)$ is convex for $\theta\in(\mu,\infty)$. We calculate
$\left(\frac{\theta}{\mu\pm\sin\theta}\right)^{\prime}=\frac{1}{\mu\pm\sin\theta}-\frac{\pm\theta\cos\theta}{(\mu\pm\sin\theta)^{2}},\quad\left(\frac{\theta}{\mu\pm\sin\theta}\right)^{\prime\prime}=\frac{\mp 2\cos\theta\pm\theta\sin\theta}{(\mu\pm\sin\theta)^{2}}+\frac{2\theta\cos^{2}\theta}{(\mu\pm\sin\theta)^{3}}.$
Since $1\geq\mu=\sin\theta_{k}>\sin\theta>0$ and $\theta>0$, we have
$\frac{2\cos\theta}{(\mu-\sin\theta)^{2}}>\frac{2\cos\theta}{(\mu+\sin\theta)^{2}},\quad\frac{2\theta\cos^{2}\theta}{(\mu-\sin\theta)^{3}}\geq\frac{\theta\sin\theta}{(\mu-\sin\theta)^{2}}.$
Hence, $\phi$ is convex for $\theta\in(\mu,\theta_{k})$. If $\mu_{+}\leq\theta_{-}$ (that is, $\Theta\cap(\mu,\theta_{k})\neq\emptyset$), then $\phi(\theta)$ for $\theta\in[\mu_{+},\theta_{-}]$ attains its maximum at either $\mu_{+}$ or $\theta_{-}$.
When $\theta_{k}<\theta<\frac{\pi}{2}$, we have
[…]
The proof can be found in Appendix A. Hence, $\phi(\theta)$ on $\theta\in[\theta_{+},\theta_{\max}]$ attains its maximum at either $\theta_{+}$ or $\theta_{\max}$ when $\theta_{+}<\theta_{\max}$ (otherwise, there is no such interval to consider). Note that if $\mu\leq C_{\mu}<\frac{1}{2}$ and $k>\pi$ then it holds that $\theta_{+}<\theta_{\max}$. Indeed, we have $h<\frac{1}{\pi}<\frac{2}{\pi+\sigma_{+}^{h}}$ so $\sin(\theta_{\max}-\frac{\sigma_{+}^{h}h}{2})=\cos(\frac{\pi+\sigma_{+}^{h}}{2}h)>1-\frac{(\pi+\sigma_{+}^{h})^{2}}{8}h^{2}>\frac{1}{2}>\sin\theta_{k}$.
Next, we shall estimate the candidates for $\max\phi(\theta)$. Using $\sin x>x-\frac{1}{3!}x^{3}$ for $x\in(0,\frac{\pi}{2})$ and $\mu-\mu_{-}=\frac{\sigma_{-}h}{2}>0$, we have
$\phi(\mu_{-})$ $=\frac{\mu_{-}-\sin\mu_{-}}{\mu-\sin\mu_{-}}\cdot\frac{\mu_{-}+\sin\mu_{-}}{\mu+\sin\mu_{-}}\cdot\frac{\mu_{-}}{\mu+\mu_{-}}\cdot\frac{1}{\mu-\mu_{-}}$
$<\frac{\mu_{-}-(\mu_{-}-\frac{1}{3!}\mu_{-}^{3})}{\mu-(\mu_{-}-\frac{1}{3!}\mu_{-}^{3})}\cdot 1\cdot\frac{1}{2}\cdot\frac{2}{\sigma_{-}h}<\frac{\mu^{3}}{3\sigma_{-}h+\mu^{3}}\cdot\frac{1}{\sigma_{-}h}=\frac{k^{3}h}{(24\sigma_{-}+k^{3}h^{2})\sigma_{-}}.$
Note that $\sigma_{-}<\pi<\frac{k}{2}$ implies $\mu<2\mu-\sigma_{-}h=2\mu_{-}$. Using also $x-\frac{1}{3!}x^{3}<\sin x<x-\frac{1}{3!}x^{3}+\frac{1}{5!}x^{5}$ for $x\in(0,\frac{\pi}{2})$ and $\mu_{-}<\mu<1$, we get
$\phi(\mu_{-})$ $>\frac{\mu_{-}-(\mu_{-}-\frac{1}{3!}\mu_{-}^{3}+\frac{1}{5!}\mu_{-}^{5})}{\mu-(\mu_{-}-\frac{1}{3!}\mu_{-}^{3}+\frac{1}{5!}\mu_{-}^{5})}\cdot\frac{\mu_{-}+\mu_{-}-\frac{1}{3!}\mu_{-}^{3}}{\mu+\mu_{-}-\frac{1}{3!}\mu_{-}^{3}}\cdot\frac{1}{3}\cdot\frac{2}{\sigma_{-}h}$
$=\frac{\mu_{-}^{3}(1-\frac{1}{20}\mu_{-}^{2})}{3\sigma_{-}h+\mu_{-}^{3}(1-\frac{1}{20}\mu_{-}^{2})}\cdot\frac{\mu_{-}+\mu_{-}-\frac{1}{3!}\mu_{-}^{3}}{\mu+\mu_{-}-\frac{1}{3!}\mu_{-}^{3}}\cdot\frac{1}{3}\cdot\frac{2}{\sigma_{-}h}$
$>\frac{\frac{\mu^{3}}{2^{3}}(1-\frac{1}{20})}{3\sigma_{-}h+\frac{\mu^{3}}{2^{3}}(1-\frac{1}{20})}\cdot\frac{\mu_{-}+\mu_{-}-\frac{1}{3!}\mu_{-}}{2\mu_{-}+\mu_{-}-\frac{1}{3!}\mu_{-}}\cdot\frac{1}{3}\cdot\frac{2}{\sigma_{-}h}>\frac{k^{3}h}{(240\sigma_{-}+k^{3}h^{2})3\sigma_{-}}.$
If $\mu_{+}<\theta_{k}$ i.e. $\Theta\cap(\mu,\theta_{k})\neq\emptyset$, then $\phi(\mu_{+})$ and $\phi(\theta_{-})$ need to be counted as candidates for the maximum. Since $\theta_{k}=\arcsin\mu$, we know $\sin\mu_{+}<\mu$, and
$\phi(\mu_{+})=\frac{\mu_{+}-\sin\mu_{+}}{\mu-\sin\mu_{+}}\cdot\frac{\mu_{+}+\sin\mu_{+}}{\mu+\sin\mu_{+}}\cdot\frac{\mu_{+}}{\mu+\mu_{+}}\cdot\frac{1}{\mu_{+}-\mu}.$
Using $1>C_{\mu}\geq\mu=\sin\theta_{k}$, $\mu_{+}\leq\theta_{-}<\theta_{k}$ and $\sigma_{-}^{h}h/2=\theta_{k}-\theta_{-}<\pi h/2$ we find
$\mu-\sin\mu_{+}$ $=\sin\theta_{k}-\sin\mu_{+}>(\theta_{k}-\mu_{+})\cos\theta_{k}\geq\left(\frac{\sigma_{-}^{h}h}{2}+\theta_{-}-\mu_{+}\right)\sqrt{1-C_{\mu}^{2}},$
$\mu-\sin\mu_{+}$ $=\sin\theta_{k}-\sin\mu_{+}<\theta_{k}-\mu_{+}<\frac{\pi h}{2}+\theta_{-}-\mu_{+}.$
Note that $\theta_{k}>\theta_{-}\geq\mu_{+}>\mu$ and $\arcsin x-x<x^{3}$ for $0<x<1$. So
$0\leq\theta_{-}-\mu_{+}<\theta_{k}-\mu=\arcsin\mu-\mu<\mu^{3}.$
Using $x-\frac{x^{3}}{3!}+\frac{x^{5}}{5!}\!>\!\sin x\!>\!x\!-\!\frac{x^{3}}{3!}$ for $x\!\in\!(0,\pi/2)$ and $\mu\!<\!\mu_{+}\!<\!2\mu<2$ (for $\mu\!>\!\pi h/2\!>\!\mu_{+}\!-\!\mu$), we have
$\frac{2\mu_{+}^{3}}{15}<\frac{\mu_{+}^{3}}{6}\left(1-\frac{\mu_{+}^{2}}{20}\right)<\mu_{+}-\sin{\mu_{+}}<\frac{\mu_{+}^{3}}{6}<\frac{4}{3}\mu^{3}.$
Combining all this, we obtain
$\frac{2}{15}\frac{(k+\sigma_{+})^{3}h}{\sigma_{+}(4\pi+k^{3}h^{2})}=\frac{2}{15}\frac{\mu_{+}^{3}}{\frac{\pi h}{2}+\mu^{3}}\cdot\frac{1}{\sigma_{+}h}<\phi(\mu_{+})<\frac{4}{3}\frac{\mu^{3}}{\frac{\sigma_{-}^{h}h}{2}\sqrt{1-C_{\mu}^{2}}}\cdot\frac{4}{\sigma_{+}h}=\frac{4}{3}\frac{k^{3}h}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}.$
If $\theta_{-}\!>\!\mu$ i.e. $\Theta\!\cap\!(\mu,\theta_{k})\!\neq\!\emptyset$, $\phi(\theta_{-})$ is a candidate for $\max\phi$. For $\mu\!=\!\sin\theta_{k}$ and $\theta_{k}\!>\!\theta_{-}$, we have
$\phi(\theta_{-})=\frac{\theta_{-}-\sin\theta_{-}}{\mu-\sin\theta_{-}}\cdot\frac{\theta_{-}+\sin\theta_{-}}{\mu+\sin\theta_{-}}\cdot\frac{\theta_{-}}{\mu+\theta_{-}}\cdot\frac{1}{\theta_{-}-\mu}.$
Using $1>C_{\mu}\geq\mu=\sin\theta_{k}$ and $\sigma_{-}^{h}h/2=\theta_{k}-\theta_{-}<\pi h/2$, we find
$\frac{\pi h}{2}>\mu-\sin\theta_{-}=\sin\theta_{k}-\sin\theta_{-}>(\theta_{k}-\theta_{-})\cos\theta_{k}\geq\frac{\sigma_{-}^{h}h}{2}\sqrt{1-C_{\mu}^{2}}.$
Using $x-\frac{x^{3}}{3!}+\frac{x^{5}}{5!}\!>\!\sin x\!>\!x\!-\!\frac{x^{3}}{3!}$ for $x\!\in\!(0,\frac{\pi}{2})$ and $\mu\!+\!\frac{h}{2}\sigma_{+}\!\leq\!\theta_{-}\!<\!\frac{\pi}{2}\mu$ (for $\theta_{-}\!<\!\theta_{k}\!<\!\frac{\pi}{2}\sin\theta_{k}$), we have
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(\mu+\frac{h}{2}\sigma_{+})^{3}}{6}<\frac{\theta_{-}^{3}}{6}\left(1-\frac{\theta_{-}^{2}}{20}\right)<\theta_{-}-\sin{\theta_{-}}<\frac{\theta_{-}^{3}}{6}<\frac{\pi^{3}}{48}\mu^{3}.$
Combining all this and using also $\theta_{-}-\mu=\theta_{-}-\mu_{+}+\mu_{+}-\mu<\mu^{3}+\frac{\sigma_{+}h}{2}$, we obtain
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{3}h}{6\pi(4\sigma_{+}+k^{3}h^{2})}<\phi(\theta_{-})<\frac{\pi^{3}}{48}\frac{\mu^{3}}{\frac{\sigma_{-}^{h}h}{2}\sqrt{1-C_{\mu}^{2}}}\cdot\frac{\pi}{2}\cdot 1\cdot\frac{2}{\sigma_{+}h}=\frac{\pi^{4}}{192}\frac{k^{3}h}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}.$
If $\theta_{+}<\theta_{\max}$, then $\theta_{+}$ needs to be considered. For $\mu\!=\!\sin\theta_{k}$ and $\theta_{k}\!<\!\theta_{+}$, we have
$\phi(\theta_{+})=\frac{\theta_{+}-\sin\theta_{+}}{\sin\theta_{+}-\mu}\cdot\frac{\theta_{+}+\sin\theta_{+}}{\mu+\sin\theta_{+}}\cdot\frac{\theta_{+}}{\mu+\theta_{+}}\cdot\frac{1}{\theta_{+}-\mu}.$
Using $1>C_{\mu}\geq\mu=\sin\theta_{k}$, $\sigma_{+}^{h}h/2=\theta_{+}-\theta_{k}<\pi h/2<\frac{kh}{4}=\frac{\mu}{2}$ and $C_{\mu}=\sin\theta_{\mu}<\cos(\sigma_{+}^{h}h/2)$ (so that $\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2}<\frac{\pi}{2}$), we find
$\frac{\pi h}{2}>\sin\theta_{+}-\mu=\sin\theta_{+}-\sin\theta_{k}>(\theta_{+}-\theta_{k})\cos\theta_{+}>\frac{\sigma_{+}^{h}h}{2}\cos\left(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2}\right).$
Using $x\!-\!\frac{x^{3}}{3!}\!+\!\frac{x^{5}}{5!}\!>\!\sin x\!>\!x\!-\!\frac{x^{3}}{3!}$ for $x\!\in\!(0,\frac{\pi}{2})$ and $\mu\!+\!\frac{\sigma_{+}^{h}h}{2}\!\leq\!\theta_{+}\!=\!\frac{\sigma_{+}^{h}h}{2}\!+\!\theta_{k}\!<\!(\frac{1}{2}+\frac{\pi}{2})\mu$, we have
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(\mu+\frac{\sigma_{+}^{h}h}{2})^{3}}{6}<\frac{\theta_{+}^{3}}{6}\left(1-\frac{\theta_{+}^{2}}{20}\right)<\theta_{+}-\sin{\theta_{+}}<\frac{\theta_{+}^{3}}{6}<\frac{(1+\pi)^{3}}{48}\mu^{3}.$
Note that $\frac{h}{2}\sigma_{+}^{h}+\frac{\mu^{3}}{6}<\theta_{+}\!-\!\mu=(\theta_{+}\!-\!\theta_{k})\!+\!(\theta_{k}\!-\!\mu)\!<\frac{h}{2}\sigma_{+}^{h}+\mu^{3}\!$. Combining all this, we get
$\phi(\theta_{+})$ $<\frac{(1+\pi)^{3}}{48}\frac{\mu^{3}}{\frac{\sigma_{+}^{h}h}{2}\cos(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2})}(\frac{1}{2}+\frac{\pi}{2})\frac{6}{3h\sigma_{+}^{h}+\mu^{3}}$
$=\frac{(1+\pi)^{4}}{8}\frac{k^{3}h}{\sigma_{+}^{h}(24\sigma_{+}^{h}+k^{3}h^{2})\cos(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2})},$
$\phi(\theta_{+})$ $>\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(\mu+\frac{\sigma_{+}^{h}h}{2})^{3}}{6\pi h(\frac{h}{2}\sigma_{+}^{h}+\mu^{3})}=\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{3}h}{6\pi(4\sigma_{+}^{h}+k^{3}h^{2})}.$
For $\mu\!=\!\sin\theta_{k}$ and $\theta_{k}\!<\!\theta_{\max}$, we have
$\phi(\theta_{\max})=\frac{\theta_{\max}^{2}-\sin^{2}\theta_{\max}}{\sin^{2}\theta_{\max}-\mu^{2}}\cdot\frac{\theta_{\max}}{\theta_{\max}^{2}-\mu^{2}}.$
Note that $\sin^{2}\theta_{\max}=\cos^{2}\frac{\pi h}{2}>(1-\frac{\pi^{2}h^{2}}{8})^{2}>1-\frac{\pi^{2}h^{2}}{4}$, $\theta_{\max}^{2}=\frac{\pi^{2}}{4}(1-2h+h^{2})$. So
$\frac{9\pi^{2}}{64}-1<\theta_{\max}^{2}-\sin^{2}\theta_{\max}<\frac{\pi^{2}}{4}-1,$
where to get the first inequality we used the assumption $0<h\leq\frac{1}{4}$. We have also
$\frac{2}{\pi}<\frac{1}{\theta_{\max}}<\frac{\theta_{\max}}{\theta_{\max}^{2}-\mu^{2}}<\frac{\pi}{2}\left(\frac{9\pi^{2}}{64}-1\right)^{-1},$
where for the second inequality the assumption $0<h\leq\frac{1}{4}$ is used. Note that
$1>\sin^{2}\theta_{\max}-\mu^{2}>\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}.$
Combining all this and using the assumption $C_{\mu}<\cos\frac{\pi h}{2}$, we get
$\frac{2}{\pi}\left(\frac{9\pi^{2}}{64}-1\right)<\phi(\theta_{\max})<\left(\frac{\pi^{2}}{4}-1\right)\left(\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}\right)^{-1}\frac{\pi}{2}\left(\frac{9\pi^{2}}{64}-1\right)^{-1}.$
We conclude by multiplying $\frac{h}{2}$ with $\phi$ to get $\psi$ (30). ∎
From Lemma 10 and Remark 8, we can see that, when $k^{3}h$ is greater than a certain constant it holds that $\psi(\xi_{\max})<\max\psi$, and the error is of order $k^{3}h^{2}$ for a fixed $f$ in the worst case; otherwise, $\max\psi=\psi(\xi_{\max})$ will be of order $h$. As discussed in Remark 6, in the case of $\max\psi=\psi(\xi_{\max})$ i.e. $k^{3}h$ less than a certain constant, we will derive an alternate estimate using the smoothness of $f$.
Before doing that, we first check whether the smoothness of $f$ improves the convergence order of $\psi(k_{-})$ (the discussion of $\psi(k_{+})$, $\psi(k_{\pm}^{h})$ would be similar). If $f\in H^{1}_{0}(0,1)$, then the estimate (29) can be modified to
$|E_{1}^{h}|_{1}^{2}=\frac{1}{2}\sum_{(\xi/\pi)=1}^{N-1}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}|\xi\hat{f}(\xi)|^{2}\leq|f_{\mathrm{low}}|_{1}^{2}\max_{\xi\in\{1,..,N-1\}\pi}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}.$ (32)
So we need only to find the maximum of
$\psi_{e}(\xi):=\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|=\left|\frac{1}{k^{2}-\xi^{2}}-\frac{1}{k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}}\right|=\frac{h^{2}}{4}\left|\frac{1}{\mu^{2}-\theta^{2}}-\frac{1}{\mu^{2}-\sin^{2}\theta}\right|=:\frac{h^{2}}{4}\phi_{e}(\theta).$ (33)
The subscript ‘e’ indicates the quantity is mainly useful for the evanescent modes with $\xi>k^{h}_{+}$, which we shall see as follows. For $\theta\in(0,\mu)$, we have
$\phi_{e}(\theta)=\frac{1}{\mu^{2}-\theta^{2}}-\frac{1}{\mu^{2}-\sin^{2}\theta},\quad\phi_{e}^{\prime}(\theta)=\frac{2\theta(\mu^{2}-\sin^{2}\theta)^{2}-(\mu^{2}-\theta^{2})^{2}\sin(2\theta)}{(\mu^{2}-\theta^{2})^{2}(\mu^{2}-\sin^{2}\theta)^{2}}>0.$
So $\max_{(0,\mu_{-}]}\phi_{e}=\phi_{e}(\mu_{-})$. The estimate of $\phi_{e}(\mu_{-})$ needs only a small modification of the estimate of $\phi(\mu_{-})$ in the proof of Lemma 10. More precisely, the term $1/(\mu+\mu_{-})$ is used now instead of $\mu_{-}/(\mu+\mu_{-})$. We already know $\mu_{-}<\mu<2\mu_{-}$. So $\frac{1}{2\mu}<1/(\mu+\mu_{-})<\frac{2}{3\mu}$ replaces $\frac{1}{3}<\mu_{-}/(\mu+\mu_{-})<\frac{1}{2}$ and gives
$\frac{k^{2}}{(240\sigma_{-}+k^{3}h^{2})\sigma_{-}}<\phi_{e}(\mu_{-})<\frac{\mu^{2}}{3\sigma_{-}h+\mu^{3}}\cdot\frac{4}{3\sigma_{-}h}=\frac{8k^{2}}{(24\sigma_{-}+k^{3}h^{2})3\sigma_{-}}.$
Hence, in terms of $\psi_{e}$ we have
$\frac{k^{2}h^{2}}{(240\sigma_{-}+k^{3}h^{2})4\sigma_{-}}<\psi_{e}(k_{-})<\frac{2k^{2}h^{2}}{(24\sigma_{-}+k^{3}h^{2})3\sigma_{-}}.$
The upper bound says that $\psi_{e}(k_{-})$ converges to zero no slower than of order $k^{2}h^{2}$. This has no difference in $h$ to the order $k^{3}h^{2}$ of $\psi(k_{-})$, but it has one order less in $k$. However, does this really mean $|E_{1}^{h}|_{1}$ converges at the same speed? Note that to make the inequality (32) sharp, or more precisely here make the estimate for $E_{1}^{h}$ restricted to $\xi\in[\pi,k_{-}]$ sharp, we need to take $f_{\mathrm{low}}$ along $\sin(k_{-}x)$ whose derivative will contribute to $|f_{\mathrm{low}}|_{1}$ with an extra factor $k$. In other words, when the frequency $k_{-}$ is active, $|E_{1}^{h}|_{1}$ is still of order $k^{3}h^{2}$, not improved by the smoothness of $f$. This is actually observed in numerical experiments. Let $f(x)=C(\sin(10\pi x)+\sin(20\pi x)+\sin(40\pi x)+\sin(80\pi x))$ with $C$ such that $\|f\|=1$, and $k\in\{10,20,40,80\}\pi+1$. For each $k$, we use a very fine mesh with $N\approx 8k^{2}$ to get a reference solution, and a sequence of doubly refined meshes starting with $N\approx k$ to compute the $H^{1}$-semi-norm of the errors. The results are shown in Figure 2 which corroborates the order $k^{3}h^{2}$. For example, at a fixed $h$, the marked points from low to high correspond to successively doubled $k$’s, and form roughly a geometric progression with the common ratio eight. Also, there are pairs of marked points, approximately at the same height but on different curves, which correspond to $k$-values in a ratio of $4$ (or $\frac{1}{4}$) and $h$-values in a ratio of $\frac{1}{8}$ (or $8$).
Figure 2: Errors in the $H^{1}$-semi-norm for $f(x)=C(\sin(10\pi x)+\sin(20\pi x)+\sin(40\pi x)+\sin(80\pi x))$.
Having understood that we would not gain more from smoothness of $f$ for the propagating modes (i.e. with $\xi\leq k_{+}^{h}$), and in the need of a better estimate when $k^{3}h$ is sufficiently small (so that $\max\psi=\psi(\xi_{\max})$ which is merely of order $h$), we derive the following alternate estimates. By the way, we show all the evanescent modes (i.e. with $\xi>k_{+}^{h}$) can not converge faster than of order $h^{2}$.
Lemma 11.
Under the assumptions of Lemma 10 and using the notation therein, it holds that
$\max_{\xi\in\{1,\ldots,N-1\}\pi}\psi_{e}=\max\{\psi_{e}(k_{-}),\psi_{e}(k_{+}),\psi_{e}(k_{-}^{h}),\psi_{e}(k_{+}^{h}),\psi_{e}(\xi_{\max})\},$
where the function $\psi_{e}$ is defined in (33). Moreover, $\psi_{e}$ is convex on $(k,\frac{\pi}{h}]$. Let $\xi_{e}$ be the unique minimal point of $\psi_{e}$ over $(k,\frac{\pi}{h}]$. Then
$\frac{k^{2}h^{2}}{(240\sigma_{-}+k^{3}h^{2})4\sigma_{-}}<\psi_{e}(k_{-})$ $\,<\frac{2k^{2}h^{2}}{(24\sigma_{-}+k^{3}h^{2})3\sigma_{-}},\,$
$\frac{1}{15}\frac{(k+\sigma_{+})^{2}h^{2}}{\sigma_{+}(4\pi+k^{3}h^{2})}<\,\psi_{e}(k_{+})$ $\,<\frac{1}{3}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}},\,$ $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{2}h^{2}}{3\pi(2+\pi)(4\sigma_{+}+k^{3}h^{2})}<\,\psi_{e}(k_{-}^{h})\,$ $<\frac{\pi^{4}}{768}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}},\,$ $\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset,$
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{2}h^{2}}{12\pi(4\sigma_{+}^{h}+k^{3}h^{2})}<\,\psi_{e}(k_{+}^{h})\,$ $<\frac{(1+\pi)^{4}k^{2}h^{2}}{32\sigma_{+}^{h}(24\sigma_{+}^{h}\!+\!k^{3}h^{2})\cos(\theta_{\mu}\!+\!\frac{\sigma_{+}^{h}h}{2})},\,$ $\text{ if }C_{\mu}<\cos\frac{\sigma_{+}^{h}h}{2},$
$\frac{(9\pi^{2}-64)h^{2}}{64\pi^{2}}<\psi_{e}(\xi_{\max})$ $\,<\frac{4(\pi^{2}-4)h^{2}}{\left(\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}\right)\left(9\pi^{2}-64\right)},\,$ $\text{ if }C_{\mu}<\cos\frac{\pi h}{2},$
$\frac{h^{2}}{18}\,<\psi_{e}(\xi_{e})$ $\,<\psi_{e}(\xi_{\max}).\,$
Remark 9.
Lemma 11 gives the better and exact order $h^{2}$ for the evanescent modes, compared to Lemma 10. It can be seen also from Lemma 11 that the propagating modes usually converge slower than the evanescent modes because the $\psi_{e}$ at the former frequencies have the extra factor $k^{2}$. We propose to use (32) and Lemma 11 only when $k^{3}h$ is sufficiently small (so that $\max\psi=\psi(\xi_{\max})$; see Lemma 10). In that case, the estimates of Lemma 11 can be further improved, but we will not make more efforts here.
Remark 10.
Lemma 11 is useful also for the $L^{2}$-norm of $E_{1}^{h}$ because
$\|E_{1}^{h}\|^{2}=\frac{1}{2}\sum_{(\xi/\pi)=1}^{N-1}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}|\hat{f}(\xi)|^{2}\leq\|f_{\mathrm{low}}\|^{2}\max_{\xi\in\{1,..,N-1\}\pi}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}.$
Hence, the exact order of $\|E_{1}^{h}\|$ is $k^{2}h^{2}$.
Proof of Lemma 11.
Recall the notation used in the proof of Lemma 10. The estimate of $\psi_{e}(k_{-})$ has been given in the above discussion, along with the result $\max_{(0,k_{-}]}\psi_{e}=\psi_{e}(k_{-})$.
When $\mu<\theta<\theta_{k}$, we have
$\phi_{e}(\theta)=\frac{1}{\theta^{2}-\mu^{2}}+\frac{1}{\mu^{2}-\sin^{2}\theta}=\frac{1}{2\mu}\left(\frac{1}{\theta-\mu}-\frac{1}{\theta+\mu}\right)+\frac{1}{2\mu}\left(\frac{1}{\mu-\sin\theta}+\frac{1}{\mu+\sin\theta}\right).$
To show that $\phi_{e}$ is convex on $(\mu,\theta_{k})$, we calculate
$2\mu\phi_{e}^{\prime\prime}(\theta)=\frac{2}{(\theta-\mu)^{3}}-\frac{2}{(\theta+\mu)^{3}}-\frac{\sin\theta}{(\mu-\sin\theta)^{2}}+\frac{\sin\theta}{(\mu+\sin\theta)^{2}}+\frac{2\cos^{2}\theta}{(\mu-\sin\theta)^{3}}+\frac{2\cos^{2}\theta}{(\mu+\sin\theta)^{3}}.$
Using $0<\sin\theta<\mu<\theta<\frac{\pi}{2}$, we find
$\frac{2}{(\theta-\mu)^{3}}>\frac{2}{(\theta+\mu)^{3}},\quad\frac{2\cos^{2}\theta}{(\mu-\sin\theta)^{3}}>\frac{\sin\theta}{(\mu-\sin\theta)^{2}}\Leftarrow\sin\theta+\sin^{2}\theta<2.$
Hence, $\phi_{e}^{\prime\prime}>0$ on $(\mu,\theta_{k})$. So the candidates for $\max\phi_{e}$ on $(\mu,\theta_{k})$ are $\phi_{e}(\mu_{+})$ and $\phi_{e}(\theta_{-})$ if $\mu_{+}<\theta_{k}$. The estimate of $\phi_{e}(\mu_{+})$ is quite similar to the estimate of $\phi(\mu_{+})$ in the proof of Lemma 10. We need only to replace $\mu_{+}/(\mu+\mu_{+})$ there with $1/(\mu+\mu_{+})$ here. Since $0<\mu<\mu_{+}$, we find
[…]
should be used. Therefore, we get
$\frac{1}{15}\frac{(k+\sigma_{+})^{2}h^{2}}{\sigma_{+}(4\pi+k^{3}h^{2})}<\,\psi_{e}(k_{+})<\frac{1}{3}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,,\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset.$
In a similar way, noting that $\mu+\theta_{-}<\mu_{+}+\theta_{k}<\mu_{+}+\frac{\pi}{2}\mu$ and using
$\frac{2}{(2+\pi)\mu_{+}}<\frac{1}{\mu+\theta_{-}}<\frac{1}{2\mu}\quad\text{ instead of }\quad\frac{1}{2}<\frac{\theta_{-}}{\mu+\theta_{-}}<1,$
we find
$\frac{(1-\frac{\pi^{2}}{80}C_{\mu}^{2})(k+\sigma_{+})^{2}h^{2}}{3\pi(2+\pi)(4\sigma_{+}+k^{3}h^{2})}<\,\psi_{e}(k_{-}^{h})<\frac{\pi^{4}}{768}\frac{k^{2}h^{2}}{\sigma_{-}^{h}\sigma_{+}\sqrt{1-C_{\mu}^{2}}}\,,\text{ if }\pi\mathbb{N}\cap(k,k^{h})\neq\emptyset.$
When $\theta_{k}<\theta<\frac{\pi}{2}$, with $\mu=\sin\theta_{k}$ we have that
[…]
The proof can be found in Appendix B. So the candidates for $\max\phi_{e}$ on $\Theta\cap(\theta_{k},\theta_{\max}]$ are $\phi_{e}(\theta_{+})$ (if $\theta_{+}<\theta_{\max}$) and $\phi_{e}(\theta_{\max})$. To estimate $\phi_{e}(\theta_{+})$, we can again modify the estimates of $\phi(\theta_{+})$ in the proof of Lemma 10. Using
[…]
we can find
$\frac{(1-\frac{(1+\pi)^{2}}{80}C_{\mu}^{2})(k+\sigma_{+}^{h})^{2}h^{2}}{12\pi(4\sigma_{+}^{h}+k^{3}h^{2})}<\,\psi_{e}(k_{+}^{h})<\frac{(1+\pi)^{4}k^{2}h^{2}}{32\sigma_{+}^{h}(24\sigma_{+}^{h}+k^{3}h^{2})\cos(\theta_{\mu}+\frac{\sigma_{+}^{h}h}{2})}\,,\text{ if }C_{\mu}<\cos\frac{\sigma_{+}^{h}h}{2}.$
For $\mu\!=\!\sin\theta_{k}$ and $\theta_{k}\!<\!\theta_{\max}$, we have
$\phi_{e}(\theta_{\max})=\frac{\theta_{\max}^{2}-\sin^{2}\theta_{\max}}{\sin^{2}\theta_{\max}-\mu^{2}}\cdot\frac{1}{\theta_{\max}^{2}-\mu^{2}}.$
Recall from the proof of Lemma 10,
$\frac{9\pi^{2}}{64}-1<\theta_{\max}^{2}-\sin^{2}\theta_{\max}<\frac{\pi^{2}}{4}-1,\quad 1>\sin^{2}\theta_{\max}-\mu^{2}>\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}.$
Since $\theta_{\max}^{2}=\frac{\pi^{2}}{4}(1-2h+h^{2})$, $0<\mu<1$ and $0<h\leq\frac{1}{4}$, we find
$\frac{\pi^{2}}{4}>\theta_{\max}^{2}-\mu^{2}>\frac{9\pi^{2}}{64}-1.$
Combining this and using the assumption $C_{\mu}<\cos\frac{\pi h}{2}$, we obtain
$\left(\frac{9\pi^{2}}{64}-1\right)\frac{4}{\pi^{2}}<\phi_{e}(\theta_{\max})<\left(\frac{\pi^{2}}{4}-1\right)\left(\cos^{2}\frac{\pi h}{2}-C_{\mu}^{2}\right)^{-1}\left(\frac{9\pi^{2}}{64}-1\right)^{-1}.$
Denote $\theta_{e}:={\xi_{e}h}/{2}$. To find the minimal point $\theta_{e}$ of $\phi_{e}$ on $\theta_{k},\frac{\pi}{2}]$, we calculate
$\phi_{e}^{\prime}(\theta)=-\frac{\sin(2\theta)}{(\sin^{2}\theta-\mu^{2})^{2}}+\frac{2\theta}{(\theta^{2}-\mu^{2})^{2}},\quad\phi_{e}^{\prime}\left(\frac{\pi}{2}\right)>0,\quad\lim_{\theta\to\theta_{k}+0}\phi_{e}^{\prime}(\theta)<0.$
Since $\phi_{e}$ is convex in $\theta_{k},\frac{\pi}{2}]$ and $\theta_{e}=\arg\min_{[\theta_{k},{\pi}/{2}]}\phi_{e}$, we have $\phi_{e}^{\prime}(\theta_{e})=0$, from which we find
$\frac{1}{\sin^{2}\theta_{e}-\mu^{2}}=\sqrt{\frac{2\theta_{e}}{\sin(2\theta_{e})}}\cdot\frac{1}{\theta_{e}^{2}-\mu^{2}}.$
Substituting this into $\phi_{e}(\theta_{e})$, and using $\frac{x}{\sin x}\!>\!1\!+\!\frac{x^{2}}{6}$ for $0\!<\!x\!<\!\pi$ and $\sqrt{1+x}\!>\!1\!+\!\frac{x}{3}$ for 0<x<30\!<\!x\!<\!3, we get
$\phi_{e}(\theta_{e})=\left(\sqrt{\frac{2\theta_{e}}{\sin(2\theta_{e})}}-1\right)\frac{1}{\theta_{e}^{2}-\mu^{2}}>\frac{4\theta_{e}^{2}}{18}\cdot\frac{1}{\theta_{e}^{2}}=\frac{2}{9}.$
Multiplying $\phi_{e}$ with $\frac{h^{2}}{4}$ gives the conclusion for $\psi_{e}$. ∎
Putting all the results on the downsampling, aliasing and operator errors together, we have an estimate of the total error in the $H^{1}$-semi-norm and $L^{2}$-norm.
Theorem 3 (Absolute error with nonzero $f$).
Suppose the problem (1) has $k>2\pi$, $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$, $f\in H_{0}^{p}(0,1)$ with $p\geq 1$, $g_{0}=g_{1}=0$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. Suppose $0<\frac{kh}{2}\leq C_{\mu}<\cos\frac{\pi h}{2}$, and there is a sequence of $h=\frac{1}{N}$ ($N\geq 4$) going to zero such that
[…]
Let $f$ be decomposed as (6). Then the finite difference scheme (9) has a unique solution $u^{h}$. Moreover,
$|u-u^{h}|_{1}\leq\left(\frac{h}{\pi^{2}-4C_{\mu}^{2}}+\frac{1}{\tilde{\sigma}_{k}}\right)\pi^{1-p}h^{p}|f_{\mathrm{high}}|_{p}+|E_{1}^{h}|_{1},$
$\|u-u^{h}\|\leq\left(\frac{h^{2}}{\pi^{2}-4C_{\mu}^{2}}+\frac{2}{\tilde{\sigma}_{k}k}\right)\pi^{-p}h^{p}|f_{\mathrm{high}}|_{p}+\|E_{1}^{h}\|,$
with $E_{1}^{h}$ given in (28), and equality is attained when $f_{\mathrm{high}}=0$. In turn, $E_{1}^{h}$ satisfies
$\frac{h^{2}}{18}|f_{\mathrm{low,e}}|_{1}\leq|E_{1}^{h}|_{1}\leq\min\{C(k,h)\|f_{\mathrm{low}}\|,C_{e}(k,h)|f_{\mathrm{low}}|_{1}\},\quad\frac{h^{2}}{18}\|f_{\mathrm{low,e}}\|\leq\|E_{1}^{h}\|\leq C_{e}(k,h)\|f_{\mathrm{low}}\|,$
where $f_{\mathrm{low,e}}:=\sum_{(\xi/\pi)=\lceil k^{h}/\pi\rceil}^{N-1}\hat{f}(\xi)\sin(\xi x)$, $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$, the upper bounds are attainable, and there exist positive constants $C_{1}$, $C_{2}$, $c$, $c_{e}$, $\tilde{C}$ and $\tilde{C}_{e}$ independent of $k$ and $h$ such that, when $k^{3}h^{2}<C_{1}$ it holds that
$ck^{3}h^{2}<C(k,h)<\tilde{C}k^{3}h^{2},\quad c_{e}k^{3}h^{2}<kC_{e}(k,h)<\tilde{C}_{e}k^{3}h^{2},$
and when $k^{3}h^{2}>C_{2}$ it holds that $C(k,h)$ or $kC_{e}(k,h)$ is greater than a positive constant.
Remark 11.
Here $C_{e}(k,h)$ is multiplied with $k$ because $|E_{1}^{h}|_{1}\leq C_{e}(k,h)|f_{\mathrm{low}}|_{1}$ holds with equality when $f_{\mathrm{low}}=\sin(\xi x)$ for some $\xi\in\pi\mathbb{N}$ satisfying $|\xi-k|<\pi$ and in this case $|f_{\mathrm{low}}|_{1}$ is of order $k$. Note that the upper bound is attained when $f$ is chosen dependently on $k$ and as a result $|f|_{p}=O(k^{p})$.
Proof.
The proof is obtained by combining (3), (7), Lemma 8, (22), (29), Lemma 10, (32) and Lemma 11. ∎
3.5 Relative error estimates of the classical 3-point centred scheme
To estimate the relative error, we recall the solution $u$ of (1) and the solution $u^{h}$ of (9) in case (i), $g_{0}=g_{1}=0$, are
$u(x)=\sum_{\xi\in\pi\mathbb{N}}\frac{\hat{f}(\xi)}{k^{2}-\xi^{2}}\sin(\xi x),\quad u^{h}(x)=\sum_{(\xi/\pi)=1}^{N-1}\frac{\widehat{f^{h}}(\xi)}{k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}}\sin(\xi x),$
$\widehat{f^{h}}(\xi)=\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi),\text{ for }\xi\in\{1,\ldots,N-1\}\pi.$
We have already shown that $f_{\mathrm{high}}=\sum_{(\xi/\pi)=N+1}^{\infty}\hat{f}(\xi)\sin(\xi x)$ causes the downsampling error (7) and aliasing error (26). Let us see how they behave in the relative sense.
Lemma 12.
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
Then, for $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$ and $p\geq 1$, the downsampling error $e_{2}^{h}$ in (3) satisfies
$\frac{\|e_{2}^{h}\|}{\|u\|}<\frac{1}{\pi^{p}}h^{p}\frac{|f_{\mathrm{high}}|_{p}}{\|f_{\mathrm{low}}\|},\quad\frac{|e_{2}^{h}|_{1}}{|u|_{1}}<\frac{1}{\pi^{p-1}}h^{p-1}\frac{|f_{\mathrm{high}}|_{p}}{|f_{\mathrm{low}}|_{1}}.$
Remark 12.
Recall from Remark 1 that as $N\to\infty$ we have $f_{\mathrm{high}}\to 0$ and $f_{\mathrm{low}}\to f$ for a fixed $f\in H^{p}_{0}(0,1)$. So the relative errors converge faster than the displayed powers of $h$. But if $|f_{\mathrm{high}}|_{p}=O(k^{p})$ and $kh=O(1)$, then Lemma 12 says the relative error is bounded in $L^{2}$ and $H^{1}$.
Remark 13.
Compared to the absolute errors of downsampling (7) and (8), the exponents of $h$ here are decreased by two. This is caused by the fact that $f_{\mathrm{low}}$ depends on $h$ and the bound for $u_{\mathrm{low}}$ becomes sharp only for increasingly oscillatory $f$ (see the following proof). If $f$ and $k$ are fixed, then $u$ is fixed and the relative error should be of the same order in $h$ as the absolute error is.
Proof of Lemma 12.
By Parseval’s identity, we have
$\frac{\|e_{2}^{h}\|^{2}}{\|u\|^{2}}=\frac{\sum_{(\xi/\pi)=N}^{\infty}\frac{1}{|k^{2}-\xi^{2}|^{2}}|\hat{f}(\xi)|^{2}}{\left(\sum_{(\xi/\pi)=1}^{N-1}+\sum_{(\xi/\pi)=N}^{\infty}\right)\frac{1}{|k^{2}-\xi^{2}|^{2}}|\hat{f}(\xi)|^{2}}=\frac{\|e_{2}^{h}\|^{2}}{\|u_{\mathrm{low}}\|^{2}+\|e_{2}^{h}\|^{2}}.$
Note that
$\|e_{2}^{h}\|\leq\max_{\xi\geq N\pi}\frac{1}{|k^{2}-\xi^{2}|\cdot|\xi|^{p}}|f_{\mathrm{high}}|_{p},\quad\|u_{\mathrm{low}}\|\geq\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}\|f_{\mathrm{low}}\|.$
We consider the case of $(N-1)^{2}\pi^{2}-k^{2}>k^{2}-\pi^{2}>0$. This is guaranteed by $\frac{2}{h}\!>\!k\!>\!\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}$:
$(N\!-\!1)^{2}\pi^{2}\!-\!k^{2}\!>\!k^{2}\!-\!\pi^{2}\Leftrightarrow((N\!-\!1)^{2}\!+\!1)\pi^{2}\!>\!2k^{2}\Leftarrow((1\!-\!h)^{2}\!+\!h^{2})\pi^{2}\!>\!8\Leftarrow h<\frac{2}{k}\!<\!\frac{\pi\!-\!\sqrt{{16}\!-\!\pi^{2}}}{2\pi}.$
In this case, we have
$\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}=\frac{1}{(N-1)^{2}\pi^{2}-k^{2}},\quad\max_{\xi\geq N\pi}\frac{1}{|k^{2}-\xi^{2}|\cdot|\xi|^{p}}=\frac{1}{N^{2}\pi^{2}-k^{2}}\frac{1}{N^{p}\pi^{p}}.$
Therefore, it holds that
$\frac{\|e_{2}^{h}\|}{\|u\|}\leq\frac{\|e_{2}^{h}\|}{\|u_{\mathrm{low}}\|}\leq\frac{(N-1)^{2}\pi^{2}-k^{2}}{N^{2}\pi^{2}-k^{2}}\frac{1}{N^{p}\pi^{p}}\frac{|f_{\mathrm{high}}|_{p}}{\|f_{\mathrm{low}}\|}<\frac{h^{p}}{\pi^{p}}\frac{|f_{\mathrm{high}}|_{p}}{\|f_{\mathrm{low}}\|}.$
For the $H^{1}$-semi-norm of $e_{2}^{h}$, we note that
$|e_{2}^{h}|_{1}\leq\max_{\xi\geq N\pi}\frac{1}{|k^{2}-\xi^{2}|\cdot|\xi|^{p-1}}|f_{\mathrm{high}}|_{p},\quad|u_{\mathrm{low}}|_{1}\geq\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}|f_{\mathrm{low}}|_{1}.$
The estimation of the relative error is similar to the previous one in the $L^{2}$-norm. ∎
Lemma 13.
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
Then, for $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$ and $p\geq 2$, the aliasing error $E_{2}^{h}$ in (27) satisfies
$\frac{\|E_{2}^{h}\|}{\|u\|}<\frac{2}{\tilde{\sigma}_{k}k\pi^{p-2}}h^{p-2}\frac{|f_{\mathrm{high}}|_{p}}{\|f_{\mathrm{low}}\|},\quad\frac{|E_{2}^{h}|_{1}}{|u|_{1}}<\frac{1}{\tilde{\sigma}_{k}\pi^{p-3}}h^{p-2}\frac{|f_{\mathrm{high}}|_{p}}{|f_{\mathrm{low}}|_{1}}.$
Proof.
We invoke Lemma 9 and Lemma 8 for bounds of $E_{2}^{h}$, and recall from the proof of Lemma 12 that
$\|u_{\mathrm{low}}\|\geq\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}\|f_{\mathrm{low}}\|,\quad|u_{\mathrm{low}}|_{1}\geq\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}|f_{\mathrm{low}}|_{1},$
$\min_{\xi\in\{1,..,N-1\}\pi}\frac{1}{|k^{2}-\xi^{2}|}=\frac{1}{(N-1)^{2}\pi^{2}-k^{2}}.$
Combining all this gives the upper bounds for $\|E_{2}^{h}\|/\|u_{\mathrm{low}}\|\!\geq\!\|E_{2}^{h}\|/\|u\|$ and $|E_{2}^{h}|_{1}/|u_{\mathrm{low}}|_{1}\!\geq\!|E_{2}^{h}|_{1}/|u|_{1}$. ∎
Now we start estimating the relative error due to operator discretization. That is, for $E_{1}^{h}$ in (27),
$\frac{\|E_{1}^{h}\|^{2}}{\|u\|^{2}}=\frac{\sum_{(\xi/\pi)=1}^{N-1}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}|\hat{f}(\xi)|^{2}}{\|u\|^{2}}=\sum_{(\xi/\pi)=1}^{N-1}w^{h}(\xi)\frac{|\lambda(\xi)-\lambda^{h}(\xi)|^{2}}{|\xi^{p}\lambda^{h}(\xi)|^{2}},$ (35)
where the weight function $w^{h}(\xi)$ is given by
$w^{h}(\xi):=\frac{|\xi^{p}\hat{u}(\xi)|^{2}}{\|u\|^{2}}=\frac{|\xi^{p}\hat{f}(\xi)|^{2}}{|\lambda(\xi)|^{2}\|u\|^{2}},\quad\text{ and }\sum_{(\xi/\pi)=1}^{N-1}w^{h}(\xi)=\frac{|u_{\mathrm{low}}|_{p}^{2}}{\|u\|^{2}}.$ (36)
Note that as $h=\frac{1}{N}\to 0$, $|u_{\mathrm{low}}|_{p}\to|u|_{p}$. It can be deduced from (35) and (36) that
$\frac{\|E_{1}^{h}\|}{\|u\|}\leq\frac{|u_{\mathrm{low}}|_{p}}{\|u\|}\max_{\xi\in\{1,..,N-1\}\pi}\left|\frac{\lambda-\lambda^{h}}{\xi^{p}\lambda^{h}}\right|,$ (37)
[…]
The relative error in the $H^{1}$-semi-norm can be analyzed in a similar way. Specifically, we have
$\frac{|E_{1}^{h}|_{1}^{2}}{|u|_{1}^{2}}=\frac{\sum_{(\xi/\pi)=1}^{N-1}\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|^{2}|\xi\hat{f}(\xi)|^{2}}{|u|_{1}^{2}}=\sum_{(\xi/\pi)=1}^{N-1}\frac{|\xi^{p+1}\hat{f}(\xi)|^{2}}{|\lambda(\xi)|^{2}|u|_{1}^{2}}\cdot\frac{|\lambda(\xi)-\lambda^{h}(\xi)|^{2}}{|\xi^{p}\lambda^{h}(\xi)|^{2}},$
which leads to
$\frac{|E_{1}^{h}|_{1}}{|u|_{1}}\leq\frac{|u_{\mathrm{low}}|_{p+1}}{|u|_{1}}\max_{\xi\in\{1,..,N-1\}\pi}\left|\frac{\lambda-\lambda^{h}}{\xi^{p}\lambda^{h}}\right|.$ (38)
So for both the $L^{2}$- and $H^{1}$-norms it is essential to find the same maximum.
Lemma 14.
Suppose $2\pi<k\not\in\pi\mathbb{N}$, $\frac{kh}{2}\leq C_{\mu}\leq\frac{3}{4}$ and a sequence of $h=\frac{1}{N}<\frac{1}{2\pi}$ satisfying
[…]
Let $\xi_{\max}:=(N-1)\pi$, $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$, $k_{\pm}^{h}\in\pi\mathbb{N}$ such that $k_{-}^{h}<k^{h}<k_{+}^{h}$ and $k_{+}^{h}-k_{-}^{h}=\pi$. Denote $\tilde{\sigma}_{\pm}^{h}:=|k-\frac{2}{h}\sin\frac{k_{\pm}^{h}h}{2}|$. Let $\lambda:=k^{2}-\xi^{2}$, $\lambda^{h}:=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$ and $\psi_{rel}(\xi):=\left|\frac{\lambda-\lambda^{h}}{\xi^{2}\lambda^{h}}\right|$. Then $\max_{\xi\in\{1,..,N-1\}\pi}\psi_{rel}(\xi)$ is attained at one of $k_{-}^{h}$, $k_{+}^{h}$ and $\xi_{\max}$ with
$\frac{1}{196\tilde{\sigma}_{-}^{h}}kh^{2}\!<\!\psi_{rel}(k_{-}^{h})\!<\!\frac{\pi^{3}}{48(2+\pi)\tilde{\sigma}_{-}^{h}}kh^{2},\quad\frac{1}{64\tilde{\sigma}_{+}^{h}}kh^{2}\!<\!\psi_{rel}(k_{+}^{h})\!<\!\frac{3\pi}{32\tilde{\sigma}_{+}^{h}}kh^{2},\quad\frac{9}{100}h^{2}\!<\!\psi_{rel}(\xi_{\max})\!<\!\frac{2}{3}h^{2},$
where $\pi\geq\tilde{\sigma}_{\pm}^{h}\geq\tilde{\sigma}_{k}$.
Proof.
Let $\mu:=\frac{kh}{2}$, $\theta:=\frac{\xi h}{2}$. We find $\psi_{rel}(\xi)=\frac{h^{2}}{4}\cdot\frac{\theta^{2}-\sin^{2}\theta}{\theta^{2}|\mu^{2}-\sin^{2}\theta|}=:\frac{h^{2}}{4}\phi_{rel}(\theta)$. When $\sin\theta<\mu$,
$\phi_{rel}(\theta)=\frac{1-\frac{\sin^{2}\theta}{\theta^{2}}}{\mu^{2}-\sin^{2}\theta}\quad\text{increases with }\theta\text{ for }0<\theta<\frac{\pi}{2}\text{ and }\sin\theta<\mu.$
So $k_{-}^{h}$ is the unique maximizer of $\psi_{rel}(\xi)$ on $(0,k_{-}^{h}]$. When $\sin\theta>\mu$, we have
$\phi_{rel}(\theta)=\frac{\theta^{2}-\sin^{2}\theta}{\theta^{2}(\sin^{2}\theta-\mu^{2})}\quad{\color[rgb]{1,0,0}\text{has no local maximum for }0<\theta_{k}:=\frac{h}{2}k^{h}=\arcsin\mu<\theta<\frac{\pi}{2}}.$
Indeed, we calculate
$\theta^{3}\left(\sin^{2}\theta-\mu^{2}\right)^{2}\phi_{rel}^{\prime}(\theta)=-2\left(\theta^{3}\sin\theta\cos\theta-\sin^{4}\theta+\mu^{2}\sin^{2}\theta-\theta\mu^{2}\sin\theta\cos\theta\right)$
and a criticial point $\theta_{0}\in(\arcsin\mu,\frac{\pi}{2})$ of $\phi_{rel}$ is a root of
$\mu^{2}=\frac{\theta^{3}\cos\theta-\sin^{3}\theta}{\theta\cos\theta-\sin\theta}=:v(\theta).$ (39)
By calculations, we find
$(\theta\cos\theta-\sin\theta)^{2}v^{\prime}(\theta)=(\theta-\sin\theta)(\theta+\sin\theta)\left(\theta\sin^{2}\theta+3\theta\cos^{2}\theta-3\sin\theta\cos\theta\right)>0,$
because $\left(\theta\sin^{2}\theta+3\theta\cos^{2}\theta-3\sin\theta\cos\theta\right)^{\prime}=4\sin^{2}\theta-4\theta\sin\theta\cos\theta>0$ for $\theta\in(0,\frac{\pi}{2})$. Hence, $v^{\prime}(\theta)>0$ and there is at most one critical point of $\phi_{rel}$. Suppose $\theta_{0}\in(\arcsin\mu,\frac{\pi}{2})$ is a critical point of $\phi_{rel}$. By substitution with (39), we find
$\theta_{0}^{5}\left(\theta_{0}^{2}-\sin^{2}\theta_{0}\right)\phi_{rel}^{\prime\prime}(\theta_{0})=2\sin\theta_{0}\cos\theta_{0}(\tan\theta_{0}-\theta_{0})\left(3\theta_{0}+\theta_{0}\tan^{2}\theta_{0}-3\tan\theta_{0}\right)>0,$
because $\left(3\theta+\theta\tan^{2}\theta-3\tan\theta\right)^{\prime}=2\tan\theta\left(\theta\sec^{2}\theta-\tan\theta\right)>0$ for $\theta\in(0,\frac{\pi}{2})$. That means $\theta_{0}$ can only be a local minimum.
It remains to estimate all the candidates $\phi_{rel}(\theta_{-})$, $\phi_{rel}(\theta_{+})$ and $\phi_{rel}(\theta_{\max})$ for $\max\phi_{rel}$, where $\theta_{\pm}:=\frac{k_{\pm}^{h}h}{2}$ and $\theta_{\max}:=\frac{(N-1)\pi h}{2}$, corresponding to $\psi_{rel}(\xi)$ at $\xi=k_{\pm}^{h},\xi_{\max}$. Note that $\frac{h}{2}\pi>\frac{h}{2}|k^{h}-k_{\pm}^{h}|=|\theta_{k}-\theta_{\pm}|\geq|\sin\theta_{k}-\sin\theta_{\pm}|=\frac{h}{2}\left|k-\frac{2}{h}\sin\frac{k_{\pm}^{h}h}{2}\right|=\frac{h}{2}\tilde{\sigma}_{\pm}^{h}\geq\frac{h}{2}\tilde{\sigma}_{k}$, $k_{-}^{h}<k^{h}=\frac{2}{h}\arcsin\frac{kh}{2}<\frac{\pi}{2}k$, $k^{h}_{-}>k^{h}-\pi>k-\pi>\frac{k}{2}$, and $k<k_{+}^{h}<k^{h}+\pi<\frac{\pi}{2}k+\frac{1}{2}k<3k$. These are useful in the following.
First, at $\theta_{-}$, we have
$\phi_{rel}(\theta_{-})=\frac{\theta_{-}^{2}-\sin^{2}\theta_{-}}{\theta_{-}^{2}(\sin^{2}\theta_{k}-\sin^{2}\theta_{-})}=\left(1-\frac{\sin\theta_{-}}{\theta_{-}}\right)\left(1+\frac{\sin\theta_{-}}{\theta_{-}}\right)\frac{1}{\sin\theta_{k}-\sin\theta_{-}}\cdot\frac{1}{\sin\theta_{k}+\sin\theta_{-}}.$
Note that $x-\frac{x^{3}}{6}<\sin x<x-\frac{x^{3}}{8}$ and $x<\frac{\pi}{2}\sin x$ for $x\in(0,\frac{\pi}{2})$. We get
$\phi_{rel}(\theta_{-})<\frac{1}{24}(k_{-}^{h})^{2}h^{2}\cdot 2\cdot\frac{2}{h\tilde{\sigma}_{-}^{h}}\cdot\frac{\pi}{(k+k_{-}^{h})h}<\frac{\pi^{3}}{12(2+\pi)\tilde{\sigma}_{-}^{h}}k,$
$\phi_{rel}(\theta_{-})>\frac{1}{32}(k_{-}^{h})^{2}h^{2}\cdot 1\cdot\frac{2}{h\tilde{\sigma}_{-}^{h}}\cdot\frac{2}{h(k+k_{-}^{h})}>\frac{1}{48\tilde{\sigma}_{-}^{h}}k.$
Second, at $\theta_{+}$, we have
$\phi_{rel}(\theta_{+})=\frac{\theta_{+}^{2}-\sin^{2}\theta_{+}}{\theta_{+}^{2}(\sin^{2}\theta_{+}-\sin^{2}\theta_{k})}=\left(1-\frac{\sin\theta_{+}}{\theta_{+}}\right)\left(1+\frac{\sin\theta_{+}}{\theta_{+}}\right)\frac{1}{\sin\theta_{+}-\sin\theta_{k}}\cdot\frac{1}{\sin\theta_{k}+\sin\theta_{+}}.$
Similar to the estimate of $\phi_{rel}(\theta_{-})$, we get
$\phi_{rel}(\theta_{+})<\frac{1}{24}(k_{+}^{h})^{2}h^{2}\cdot 2\cdot\frac{2}{h\tilde{\sigma}_{+}^{h}}\cdot\frac{\pi}{(k+k_{+}^{h})h}<\frac{3\pi}{8\tilde{\sigma}_{+}^{h}}k,$
$\phi_{rel}(\theta_{+})>\frac{1}{32}(k_{+}^{h})^{2}h^{2}\cdot 1\cdot\frac{2}{h\tilde{\sigma}_{+}^{h}}\cdot\frac{2}{h(k+k_{+}^{h})}>\frac{1}{16\tilde{\sigma}_{+}^{h}}k.$
Finally, at $\theta_{\max}=\frac{\pi}{2}(1-h)$, we have
$\phi_{rel}(\theta_{\max})=\frac{\theta_{\max}^{2}-\sin^{2}\theta_{\max}}{\theta_{\max}^{2}(\sin^{2}\theta_{\max}-\sin^{2}\theta_{k})}=\frac{\frac{\pi^{2}}{4}(1-h)^{2}-\cos^{2}\frac{\pi h}{2}}{\frac{\pi^{2}}{4}(1-h)^{2}(\cos^{2}\frac{\pi h}{2}-\frac{k^{2}h^{2}}{4})}.$
From the assumption $0<h<\frac{1}{2\pi}$, using also $\cos x>1-\frac{1}{2}x^{2}$ for $x\in(0,\frac{\pi}{2})$ we have $\cos^{2}\frac{\pi h}{2}=\frac{1}{2}(1+\cos(\pi h))>1-\frac{\pi^{2}h^{2}}{4}>\frac{15}{16}$. Combining this with the assumption $\frac{kh}{2}<\frac{3}{4}$ gives
$\frac{9}{25}=1-\frac{4}{\pi^{2}(\frac{3}{\pi}-\frac{1}{2\pi})^{2}}<\sec^{2}\frac{\pi h}{2}-\frac{4}{\pi^{2}(1-h)^{2}}<\phi_{rel}(\theta_{\max})<\frac{1}{\cos^{2}\frac{\pi h}{2}-\frac{k^{2}h^{2}}{4}}<\frac{8}{3}.$
To conclude for $\psi_{rel}$, we need only to multiply $\phi_{rel}$ with $\frac{h^{2}}{4}$. ∎
Now we can summarize the relative error.
Theorem 4 (Relative error with nonzero $f$).
Suppose the problem (1) has $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$, $f\in H_{0}^{p}(0,1)$ with $p\geq 2$, $g_{0}=g_{1}=0$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. Suppose $0<\frac{kh}{2}\leq C_{\mu}\leq\frac{3}{4}$, and there is a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$, $k_{\pm}^{h}\in\pi\mathbb{N}$ such that $k_{-}^{h}<k^{h}<k_{+}^{h}$ and $k_{+}^{h}-k_{-}^{h}=\pi$. Let $f$ be decomposed as in (6). Then the finite difference scheme (9) has a unique solution $u^{h}$. Moreover,
$\frac{\|u-u^{h}\|}{\|u\|}\leq\left(\frac{1}{\pi^{p}}h^{2}+\frac{2}{\tilde{\sigma}_{k}k\pi^{p-2}}\right)h^{p-2}\frac{|f_{\mathrm{high}}|_{p}}{\|f_{\mathrm{low}}\|}+\frac{\|E_{1}^{h}\|}{\|u\|},$
$\frac{|u-u^{h}|_{1}}{|u|_{1}}\leq\left(\frac{1}{\pi^{p-1}}h+\frac{1}{\tilde{\sigma}_{k}\pi^{p-2}}\right)h^{p-2}\frac{|f_{\mathrm{high}}|_{p}}{|f_{\mathrm{low}}|_{1}}+\frac{|E_{1}^{h}|_{1}}{|u|_{1}},$
with $E_{1}^{h}$ given in (28), and equality is attained when $f_{\mathrm{high}}=0$. In turn, $E_{1}^{h}$ satisfies
$\frac{\|E_{1}^{h}\|}{\|u\|}\leq\frac{|u_{\mathrm{low}}|_{2}}{\|u\|}C_{rel}(k,h),\quad\frac{|E_{1}^{h}|_{1}}{|u|_{1}}\leq\frac{|u_{\mathrm{low}}|_{3}}{|u|_{1}}C_{rel}(k,h),$
where the upper bounds are attained when $\hat{f}(\xi)=0$ for all $\xi\in\{1,\ldots,N-1\}\pi$ but one of $k_{\pm}^{h}$, and there exist positive constants $c_{rel}$ and $\tilde{C}_{rel}$ independent of $k$ and $h$ such that
$c_{rel}kh^{2}<C_{rel}(k,h)<\tilde{C}_{rel}kh^{2}.$
In particular, when the upper bounds for $E_{1}^{h}$ relative to $u$ are attained, e.g. $f(x)=2\sin(k_{\pm}^{h}x)$, we have
$u=\frac{f}{k^{2}-(k_{\pm}^{h})^{2}},\quad\|u\|=\frac{\|f\|}{k^{2}-(k_{\pm}^{h})^{2}},\quad|u|_{p}=|u_{\mathrm{low}}|_{p}=\frac{\|f\|(k_{\pm}^{h})^{p}}{k^{2}-(k_{\pm}^{h})^{2}},$
and hence the exact order of the relative errors are $k^{3}h^{2}$, or more precisely
$\frac{\|E_{1}^{h}\|}{\|u\|}=\frac{|E_{1}^{h}|_{1}}{|u|_{1}}=(k_{\pm}^{h})^{2}C_{rel}(k,h)\in(k_{\pm}^{h})^{2}kh^{2}(c_{rel},\tilde{C}_{rel}).$
Proof.
The proof is obtained by combining Lemma 12, Lemma 13, (37) (38) and Lemma 14. ∎
4 Visual analysis of dispersion correction schemes
We have already seen from (2), (10) and Section 3.1 that for the zero source problem the discretization error is essentially in the discrete wavenumber $k^{h}$. This becomes the motivation of dispersion correction. In 1D, $k^{h}$ is simply a constant depending on $kh$ for a given linear scheme, and the error in $k^{h}$ can be completely removed; see e.g. [5, 14].
For the classical 3-point centred scheme (9), one possibility adopted in [14] is to modify the wavenumber $k$ to be used in the scheme (9) and get the new scheme
$(\tilde{k}^{2}-\frac{2}{h^{2}})u^{h}(x_{j})+\frac{1}{h^{2}}u^{h}(x_{j-1})+\frac{1}{h^{2}}u^{h}(x_{j+1})=f(x_{j}),\quad u^{h}(x_{0})=g_{0},\quad u^{h}(x_{N})=g_{1}.$ (40)
When $f=0$, the solution of (40) is given by (10) which is copied here,
$u^{h}(x_{j})=\frac{g_{1}-g_{0}\cos k^{h}}{\sin k^{h}}\sin(k^{h}x_{j})+g_{0}\cos(k^{h}x_{j}),\quad x_{j}=jh,\;j=0,1,..,N,$
but with $k^{h}:=\frac{2}{h}\arcsin\frac{\tilde{k}h}{2}$ ($\tilde{k}$ replacing $k$). The idea is to let $k^{h}=k$ by well choosing $\tilde{k}$. That is, solve
$\frac{2}{h}\arcsin\frac{\tilde{k}h}{2}=k\quad\text{ to find }\quad\tilde{k}=\frac{2}{h}\sin\frac{kh}{2}.$
Then, the discretization error of the new scheme vanishes when $f=0$. An error analysis when $f\neq 0$ is given by [9] using the framework “consistency + stability $\Rightarrow$ convergence”, which shows the maximum error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor.
Another possibility adopted by [5] is to scale the discrete Laplacian but retain the original wavenumber $k$. This amounts to multiply the left hand side (only) of (40) with $\frac{k^{2}}{\tilde{k}^{2}}$ and get
$\left(k^{2}-\frac{k^{2}}{\tilde{k}^{2}}\frac{2}{h^{2}}\right)u^{h}(x_{j})+\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j-1})+\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j+1})=f(x_{j}),\quad u^{h}(x_{0})=g_{0},\quad u^{h}(x_{N})=g_{1}.$ (41)
When $f=0$, the scheme (41) is equivalent to (40). The original work of [5] is in the finite element framework with the source treated by integral. The idea is to “eliminate the phase lag $k-k^{h}$ for as many right-hand sides as possible”, not only for $f=0$ but also for piecewise constant $f$. For a different problem than (1) with the right boundary condition being $u^{\prime}-\mathrm{i}\,ku=0$, it was shown in [5] that the $H^{1}$-semi-norm of the error is bounded from above by $h\|f\|_{H^{1}}$ up to a constant factor.
Remark 14.
If we multiply $f$ (only) in (40) with $\frac{\tilde{k}^{2}}{k^{2}}$, then we obtain a scheme equivalent to (41).
As we have seen, the 1D dispersion correction is determined up to a constant factor. If we follow a third approach proposed by [40] originally for 2D and 3D problems, then we would get
$\left(k^{2}\!-\!\frac{k^{2}}{\tilde{k}^{2}}\frac{2}{h^{2}}\right)u^{h}(x_{j})\!+\!\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j-1})\!+\!\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j+1})\!=\!\frac{kh}{2}\cot\frac{kh}{2}f(x_{j}),\;u^{h}(x_{0})\!=\!g_{0},\;u^{h}(x_{N})\!=\!g_{1}.$ (42)
To evaluate the HZthree dispersion free schemes (40), (41) and (42), we use the Fourier analysis as a visual tool. Since the downsampling and aliasing errors are smaller (for smooth $f$) than the operator error, we will focus on the latter. The operator symbols of (40), (41) and (42), that map $u^{h}$ to $f^{h}$, are
$\lambda^{h}_{k}\!:=\!\tilde{k}^{2}\!-\!\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\;\;\lambda_{\Delta}^{h}\!:=\!k^{2}\!-\!\frac{k^{2}}{\tilde{k}^{2}}\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\;\;{\color[rgb]{1,0,0}\lambda_{\Delta,f}^{h}\!:=\!\frac{2}{kh}\tan\frac{kh}{2}\lambda_{\Delta}^{h},}\;\;\xi\!\in\!\{1,\ldots,N\!-\!1\}\pi\text{ for }h\!=\!\frac{1}{N}.$ (43)
Recall that the continuous symbol is $\lambda=k^{2}-\xi^{2}$, and the symbol errors are defined as
$\psi:=\left|\frac{\xi}{\lambda}-\frac{\xi}{\lambda^{h}}\right|,\quad\psi_{e}:=\left|\frac{1}{\lambda}-\frac{1}{\lambda^{h}}\right|,\quad\psi_{rel}:=\left|\frac{\lambda-\lambda^{h}}{\xi^{2}\lambda^{h}}\right|,\quad\xi\in\{1,\ldots,N-1\}\pi\text{ for }h=\frac{1}{N},$ (44)
where $\lambda^{h}$ is the discrete symbol. For the classical centred scheme, $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$. For the scheme (40), (41) or (42), it is $\lambda^{h}_{k}$, $\lambda_{\Delta}^{h}$ or $\lambda_{\Delta,f}^{h}$ in (43). From (29), (32), (37) and (38), it is essentially the symbol error $\psi$, $\psi_{e}$ or $\psi_{rel}$ that determines the convergence in terms of absolute/relative error in the $L^{2}$- or $H^{1}$-semi-norm.
Figure 3 shows the symbol errors $\psi_{e}$ (44) for the four schemes under $h$-refinement. We see that the dispersion free schemes have lower maxima of $\psi_{e}$ than the classical one, at the price of larger $\psi_{e}$ at smaller $\xi$ (roughly $\frac{5}{6}k$ or $\frac{2}{3}k$). We also see that the scheme (42) modifying both the discrete Laplacian and the source performs the best. The order of $\max\psi_{e}$ is $h^{2}$, which corroborates Lemma 11.
Figure 3: Symbol errors $\psi_{e}$ (44) for the classical (9) & dispersion free (40), (41), (42): $h$-refinement
Next, the influence of the wavenumber $k$ is demonstrated, see Figure 4 in which $k$ is quadrupled while $kh$ is halved. This clearly shows that the classical scheme can not converge in this setting with $k^{3}h^{2}$ fixed. This corresponds to the order $k^{3}h^{2}$ proved in Lemma 10. Figure 4 shows also that the schemes (40) and (41) have the symbol error of order $k^{2}h^{2}$ which translates to the same order of $|u-u^{h}|_{1}$ by (29), while the scheme (42) is even better and of  [^2] This is because the order in $h$ for fixed $k$ must be 2, and it is observed that the scaling $k\to 4k$, $h\to h/8$ makes the error scaled by $4^{3}/8^{3}=2^{-3}$, so that $4^{\alpha}/8^{2}=2^{-3}$ gives the right exponent $\alpha=\frac{3}{2}$ for $k$. order $k^{3/2}h^{2}$.
Figure 4: Symbol errors $\psi$ (44) for the classical (9) & dispersion free (40), (41), (42): khkh-refinement
Finally, the relative symbol error $\psi_{rel}$ is visualized in Figure 5 where $kh$ is fixed while $k$ is doubled. According to (38), $\max\psi_{rel}$ has to be multiplied with $|u_{\mathrm{low}}|_{3}/|u|_{1}$ to give the upper bound of $|u-u^{h}|_{1}/|u|_{1}$, and when the upper bound is attained, $|u_{\mathrm{low}}|_{3}/|u|_{1}$ may contribute some powers of $k$ (see Theorem 4). For the classical scheme, we see from Figure 5 that $\psi_{rel}$ is of order $kh^{2}$ which corroborates Lemma 14. For the dispersion free schemes (40) and (42), the relative symbol error $\psi_{rel}$ does not decrease with $h$ (while $kh$ is fixed) at the first frequency $\xi=\pi$, which slows down also the convergence at the other low frequencies. The dispersion free scheme (41) modifying the discrete Laplacian has an almost equidistributed $\psi_{rel}$ over $\xi$ as seen in the figure, and $\max\psi_{rel}$ is attained at $\xi=(N-1)\pi$ (order $k$) and is of order $h^{2}$. So the relative error $|u-u^{h}|_{1}/|u|_{1}$ of the scheme (41) is of order $k^{2}h^{2}$.
Figure 5: Symbol errors $\psi_{rel}$ (44) for the classical (9) & dispersion free (40), (41), (42): khkh fixed
5 Discussions
There are some limitations of the proposed approach in this work. First, it only works on uniform grids of simple geometry. Second, with radiation boundary conditions the generalized Fourier decomposition based on the Sturm-Liouville problem will lead to a non-orthogonal basis, and the current approach does not work directly. It might be interesting to try the singular value decomposition i.e. considering the least-squared form of the Sturm-Liouville problem, which diagonalises the Helmholtz operator with two orthogonal bases.
However, the proposed approach provides a new tool for quantitatively evaluating linear schemes, in complement to dispersion analysis. The latter considers only the zero source case and does not lead to error estimates.
For future work, we will consider models in 2D/3D with simple geometries and analyze some compact high-order schemes, which is more challenging e.g. in finding the maximum of a multivariable symbol. Moreover, it may be worthwhile to try optimizing discretization parameters based on the symbol errors.
It is still an open problem to give wavenumber explicit error estimates of finite difference schemes actually used in applications for variable coefficient Helmholtz problems. The variational theory of FDMs developed in [28] may be useful.
Acknowledgement
We thank Dr Haoran Chen from XJTLU and River Li from math.stackexchange.com for helpful discussions on some analysis.
Appendix A Proof of the claim (31)
The claim (31) is
[…]
where $\theta_{k}:=\arcsin\mu$. The following is a proof.
Proof.
We calculate
$\phi^{\prime}(\theta)=\frac{2\theta^{2}}{\left(\theta^{2}-\mu^{2}\right)^{2}}-\frac{1}{\theta^{2}-\mu^{2}}+\frac{1}{\sin^{2}\theta-\mu^{2}}-\frac{2\theta\sin\theta\cos\theta}{\left(\sin^{2}\theta-\mu^{2}\right)^{2}}=:\frac{\mathcal{P}(\theta)}{\left(\sin^{2}\theta-\mu^{2}\right)^{2}\left(\theta^{2}-\mu^{2}\right)^{2}}.$
Let $\nu=\mu^{2}$. Then a critical point $\theta_{*}$ of $\phi(\theta)$ is the root of
$\mathcal{P}(\theta)=$ $-2\theta^{5}\sin\theta\cos\theta+\theta^{4}\sin^{2}\theta+\theta^{2}\sin^{4}\theta+\left(-\theta^{4}+4\theta^{3}\sin\theta\cos\theta-4\theta^{2}\sin^{2}\theta+\sin^{4}\theta\right)\nu+$
$\left(3\theta^{2}-\sin^{2}\theta-2\theta\sin\theta\cos\theta\right)\nu^{2}=0.$
The above equation is quadratic in $\nu$. By calculation, we find
$\nu=\frac{\theta^{4}-\sin^{4}\theta-4\theta^{3}\sin\theta\cos\theta+4\theta^{2}\sin^{2}\theta\pm\frac{1}{4\sqrt{2}}{(2\theta^{2}+\cos(2\theta)-1)}\sqrt{t}}{2\left(3\theta^{2}-\sin\theta(\sin\theta+2\theta\cos\theta)\right)}=:w_{\pm}(\theta),$
where
$t=128\theta^{3}\sin\theta\cos\theta+8(\theta^{2}-\sin^{2}\theta)^{2}.$
We calculate
$w_{+}(\theta)-\sin^{2}\theta=\frac{1}{8}\frac{(\theta^{2}-\sin^{2}\theta)\left(4\theta^{2}-8\theta\sin(2\theta)+2\cos(2\theta)-2+\sqrt{2t}\right)}{3\theta^{2}-\sin\theta(\sin\theta+2\theta\cos\theta)}.$
The second factor in the numerator is actually positive, which makes $w_{+}(\theta)>\sin^{2}\theta$. Indeed, we have
$2t-\left(4\theta^{2}-8\theta\sin(2\theta)+2\cos(2\theta)-2\right)^{2}=64\theta\sin\theta\cos\theta\left(6\theta^{2}-2\theta\sin(2\theta)+\cos(2\theta)-1\right)$
and
$\left(6\theta^{2}-2\theta\sin(2\theta)+\cos(2\theta)-1\right)^{\prime}=12\theta-4\sin(2\theta)-4\theta\cos(2\theta)>0\text{ for }0<\theta<\frac{\pi}{2}.$
So a critical point of $\phi$ must satisfy $\nu=w_{-}(\theta)$. We are going to show that $w_{-}^{\prime}(\theta)>0$ for $0<\theta<\frac{\pi}{2}$. By calculation, we get
$w_{-}^{\prime}(\theta)=\frac{16\theta v(\theta)}{8\left(3\theta^{2}-\sin^{2}\theta-2\theta\sin\theta\cos\theta\right)^{2}\sqrt{t}}\quad\text{ with }$
[…]
$2\sqrt{2}\sin^{8}\theta-4\sqrt{2}\sin^{6}\theta-4\sqrt{2}\theta\sin^{5}\theta\cos\theta-\theta^{4}\sqrt{t}+5\theta^{4}\sqrt{t}\sin^{2}\theta+2\theta^{3}\sqrt{t}\sin\theta\cos\theta$-
$4\theta^{2}\sqrt{t}\sin^{4}\theta-\sqrt{t}\sin^{6}\theta+\sqrt{t}\sin^{4}\theta-2\theta\sqrt{t}\sin^{3}\theta\cos\theta.$
For $0<\theta\leq 1$, we claim
$a(\theta):=8\sqrt{2}\theta^{2}-\frac{8\sqrt{2}\theta^{4}}{3}<\sqrt{t}<\frac{7\theta^{6}}{30\sqrt{2}}-\frac{8\sqrt{2}\theta^{4}}{3}+8\sqrt{2}\theta^{2}=:b(\theta).$ (45)
Indeed, $t=64\theta^{3}\sin(2\theta)+8(\theta-\sin(\theta))^{2}(\theta+\sin(\theta))^{2}$ can be enlarged using
$\theta-\sin(\theta)$ $<\frac{\theta^{3}}{6}-\frac{\theta^{5}}{120}+\frac{\theta^{7}}{5040}$
$\theta+\sin(\theta)$ $<2\theta-\frac{\theta^{3}}{6}+\frac{\theta^{5}}{120}$
$\sin(2\theta)$ $<2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{8\theta^{7}}{315}+\frac{4\theta^{9}}{2835}$
for $0<\theta<\frac{\pi}{2}$. Denote the enlarged $t$ as $\tilde{t}$. Then we have
$b(\theta)^{2}-\tilde{t}$= $\left(\frac{584\theta^{10}}{945}-\frac{2173\theta^{12}}{22680}\right)+\left(\frac{53\theta^{14}}{18900}-\frac{1933\theta^{16}}{11907000}\right)+\left(\frac{19\theta^{18}}{2976750}-\frac{1921\theta^{20}}{11430720000}\right)+$
$\left(\frac{31\theta^{22}}{11430720000}-\frac{\theta^{24}}{45722880000}\right)>0\text{ for }0<\theta<1.$
Hence, $\sqrt{t}<b(\theta)$. We can shrink $t$ using
$\theta-\sin(\theta)$ $>\frac{\theta^{3}}{6}-\frac{\theta^{5}}{120}$
$\theta+\sin(\theta)$ $>2\theta-\frac{\theta^{3}}{6}$
$\sin(2\theta)$ $>2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{8\theta^{7}}{315}$
for $0<\theta<\frac{\pi}{2}$. Denote the shrunk $t$ as $\underset{\sim}{t}$. Then we have
$\underset{\sim}{t}-a(\theta)^{2}=\left(\frac{56\theta^{8}}{15}-\frac{352\theta^{10}}{189}\right)+\left(\frac{47\theta^{12}}{2025}-\frac{2\theta^{14}}{2025}\right)+\frac{\theta^{16}}{64800}>0\text{ for }0<\theta<1.$
Substituting (45) and
$\theta-\frac{\theta^{3}}{6}<\sin\theta<\theta-\frac{\theta^{3}}{6}+\frac{\theta^{5}}{120}$
$1-\frac{\theta^{2}}{2}+\frac{\theta^{4}}{24}-\frac{\theta^{6}}{720}<\cos\theta<1-\frac{\theta^{2}}{2}+\frac{\theta^{4}}{24}$
into $v(\theta)$ with upper bounds into negative terms and lower bounds into positive terms, yields
$v(\theta)>$ $\left(\frac{389\theta^{10}}{90\sqrt{2}}-\frac{1871\theta^{12}}{270\sqrt{2}}\right)+\left(\frac{17071\theta^{14}}{5400\sqrt{2}}-\frac{914\sqrt{2}\theta^{16}}{2025}\right)+\left(\frac{575611\theta^{18}}{2916000\sqrt{2}}-\frac{4573\theta^{20}}{144000\sqrt{2}}\right)+$
$\left(\frac{140789\theta^{22}}{38880000\sqrt{2}}-\frac{1263749\theta^{24}}{4199040000\sqrt{2}}\right)+\left(\frac{58373\theta^{26}}{3110400000\sqrt{2}}-\frac{49411\theta^{28}}{55987200000\sqrt{2}}\right)+$
$\left(\frac{17221\theta^{30}}{559872000000\sqrt{2}}-\frac{139\theta^{32}}{186624000000\sqrt{2}}\right)+\left(\frac{\theta^{34}}{89579520000\sqrt{2}}-\frac{7\theta^{36}}{89579520000000\sqrt{2}}\right)>0$
for $0<\theta<\sqrt{\frac{1167}{1871}}$. Note that $\sqrt{\frac{1167}{1871}}>\frac{7}{10}$. For $\frac{7}{10}\leq\theta\leq\frac{\pi}{2}$, we make use of interval arithmetic in Mathematica to show $v(\theta)>0$. The code is given below.
a = 7/10; b = Pi/2; n = 100; (*Number of subdivisions*) subintervals = Subdivide[a, b, n];
allIntervalsPositive = True; (*Check each subinterval*)
For[i = 1, i <= n, i++,
    smallInt = Interval[{subintervals[[i]], subintervals[[i + 1]]}];
    vSmallInt = v[smallInt];
    If[Min[vSmallInt] <= 0,  allIntervalsPositive = False;  Break[];];];
If[allIntervalsPositive, Print["v(x) > 0 in [", a, ", ", b, "]"],
    Print["Potential root found in subinterval: ", smallInt]]
In summary, $w_{-}^{\prime}(\theta)>0$ on $\theta\in(0,\frac{\pi}{2})$. Therefore, in $\theta\in(0,\frac{\pi}{2})$, $\nu=w_{-}(\theta)$ has at most one root, and $\phi(\theta)$ has at most one critical point. Note that $\mathcal{P}(\theta=\arcsin\mu)=-2\theta(\theta-\sin\theta)^{2}\sin\theta(\theta+\sin\theta)^{2}\cos\theta<0$ for $0<\mu<1$, and $\mathcal{P}(\theta=\frac{\pi}{2})=-\frac{1}{16}(\mu^{2}-1)\left((16-12\pi^{2})\mu^{2}+\pi^{4}+4\pi^{2}\right)>0$ for $0<\mu<1$. So $\phi(\theta)$ in $\theta_{k}<\theta<\frac{\pi}{2}$ is decreasing before the only critical point and then increasing after. That means $\phi(\theta)$ has no local maximum in $\theta_{k}<\theta<\frac{\pi}{2}$. ∎
Appendix B Proof of the claim (34)
We fist copy the claim (34) here:
[…]
where $\theta_{k}:=\arcsin\mu$. Then we give the proof as follows.
Proof.
We calculate
$\phi_{e}^{\prime}(\theta)=\frac{2\theta}{\left(\theta^{2}-\mu^{2}\right)^{2}}-\frac{2\sin\theta\cos\theta}{\left(\sin^{2}\theta-\mu^{2}\right)^{2}}=:2\frac{\mathcal{P}(\theta)}{\left(\theta^{2}-\mu^{2}\right)^{2}\left(\mu^{2}-\sin^{2}\theta\right)^{2}}.$
Let $\nu=\mu^{2}$. Then a critical point $\theta_{*}$ of $\phi_{e}(\theta)$ is the root of
$\mathcal{P}(\theta)=-\theta^{4}\sin\theta\cos\theta+\theta\sin^{4}\theta+\left(2\theta^{2}\sin\theta\cos\theta-2\theta\sin^{2}\theta\right)\nu+(\theta-\sin\theta\cos\theta)\nu^{2}=0$
By solving for $\nu$ from the above equation, we find $\theta_{*}$ must satisfy one of the following equations
$\nu=\frac{\theta\sin^{2}\theta-\theta^{2}\sin\theta\cos\theta\pm\sqrt{\theta}\sqrt{\sin\theta\cos\theta}\left(\theta^{2}-\sin^{2}\theta\right)}{\theta-\sin\theta\cos\theta}=:w_{\pm}(\theta).$
We calculate and use $x>\sin x>0$ for $0<x<\pi$ to find
$w_{+}(\theta)-\sin^{2}\theta=2\frac{\theta^{2}-\sin^{2}\theta}{\theta-\sin\theta\cos\theta}\left(\sqrt{2\theta}\sqrt{\sin(2\theta)}-\sin(2\theta)\right)>0.$
So the only equation that $\theta_{*}$ needs to satisfy is $\nu=w_{-}(\theta)$. We claim that $w_{-}^{\prime}(\theta)>0$ and thus $w_{-}(\theta)$ is increasing for $0<\theta<\frac{\pi}{2}$. This will be shown for $0<\theta\leq 1$ and $1\leq\theta<\frac{\pi}{2}$ separately. We calculate
$\left(2\sqrt{\theta}\sqrt{\sin\theta\cos\theta}(\sin\theta\cos\theta-\theta)^{2}\right)w_{-}^{\prime}(\theta)$
$=$ $2\theta^{3/2}\sin^{2}\theta\cos^{2}\theta\sqrt{\sin\theta\cos\theta}-2\theta^{7/2}\cos^{2}\theta\sqrt{\sin\theta\cos\theta}-2\theta^{3/2}\sin^{4}\theta\sqrt{\sin\theta\cos\theta}+$
$2\theta^{7/2}\sin^{2}\theta\sqrt{\sin\theta\cos\theta}+2\theta^{5/2}\sin\theta\cos\theta\sqrt{\sin\theta\cos\theta}+\theta^{4}\sin^{2}\theta-\theta^{4}\cos^{2}\theta-\theta^{3}\sin\theta\cos^{3}\theta+$
[…]
$\theta\sin^{5}\theta\cos\theta-\theta\sin^{3}\theta\cos\theta-2\sqrt{\theta}\sin^{3}\theta\cos\theta\sqrt{\sin\theta\cos\theta}=:v(\theta).$
We note that $v(\theta)$ is a sum of positive terms and negative terms.
For $0<\theta\leq 1$, we first show that
$\sqrt{\theta}-\frac{\theta^{5/2}}{3}+\frac{\theta^{9/2}}{90}-\frac{\theta^{13/2}}{180}<\sqrt{\sin\theta\cos\theta}<\sqrt{\theta}-\frac{\theta^{5/2}}{3}+\frac{\theta^{9/2}}{90},$ (46)
or equivalently
$a(\theta):=2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{\theta^{7}}{27}+\frac{31\theta^{9}}{4050}-\frac{\theta^{11}}{4050}+\frac{\theta^{13}}{16200}<\sin(2\theta)<2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{2\theta^{7}}{135}+\frac{\theta^{9}}{4050}=:b(\theta).$
By Maclaurin series, we can show for $0<\theta\leq 1$ that
$c(\theta):=2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{8\theta^{7}}{315}+\frac{4\theta^{9}}{2835}-\frac{8\theta^{11}}{155925}<\sin(2\theta)<2\theta-\frac{4\theta^{3}}{3}+\frac{4\theta^{5}}{15}-\frac{8\theta^{7}}{315}+\frac{4\theta^{9}}{2835}=:d(\theta)$
We note for $0<\theta\leq 1$ that
$c(\theta)-a(\theta)=\left(\frac{11\theta^{7}}{945}-\frac{59\theta^{9}}{9450}\right)+\left(\frac{61\theta^{11}}{311850}-\frac{\theta^{13}}{16200}\right)>0,\quad b(\theta)-d(\theta)=\frac{2\theta^{7}}{189}-\frac{11\theta^{9}}{9450}>0.$
We substitute (46) and
$\theta-\frac{\theta^{3}}{6}+\frac{\theta^{5}}{120}-\frac{\theta^{7}}{5040}<\sin\theta<\theta-\frac{\theta^{3}}{6}+\frac{\theta^{5}}{120}$
$1-\frac{\theta^{2}}{2}+\frac{\theta^{4}}{24}-\frac{\theta^{6}}{720}<\cos\theta<1-\frac{\theta^{2}}{2}+\frac{\theta^{4}}{24}$
into $v(\theta)$ with upper bounds into negative terms and lower bounds into positive terms, and get
$v(\theta)>$ $\left(\frac{92\theta^{10}}{315}-\frac{7619\theta^{12}}{37800}\right)+\left(\frac{5207\theta^{14}}{85050}-\frac{678781\theta^{16}}{63504000}\right)+\left(\frac{267361\theta^{18}}{228614400}-\frac{9557\theta^{20}}{120960000}\right)+\left(\frac{11827\theta^{22}}{4572288000}\right)+$
$\left(\frac{299807\theta^{24}}{5761082880000}-\frac{2240017\theta^{26}}{230443315200000}\right)+\left(\frac{78433\theta^{28}}{172832486400000}-\frac{57977\theta^{30}}{5530639564800000}\right)+$
$\left(\frac{73\theta^{32}}{592568524800000}-\frac{\theta^{34}}{1185137049600000}\right)>0\quad\text{for }0<\theta\leq 1.$
Therefore, $w_{-}^{\prime}(\theta)>0$ for $0<\theta\leq 1$.
For $1\leq\theta<\frac{\pi}{2}$, we make use of interval arithmetic in Mathematica. The code is similar to that in Appendix A. Running the code, we get the result v(x)>0 in [1,Pi/2]. Hence, $w_{-}^{\prime}(\theta)>0$ for $1\leq\theta<\frac{\pi}{2}$.
Now we have proved $w_{-}$ is increasing in $\theta\in(0,\frac{\pi}{2})$. So $\nu=w_{-}(\theta)$ has at most one root in $\theta\in(0,\frac{\pi}{2})$. That is, $\phi_{e}(\theta)$ has at most one critical point in $\theta\in(0,\frac{\pi}{2})$. Note that
$\mathcal{P}(\theta=\arcsin\mu)=-(\theta-\sin\theta)^{2}\sin\theta(\theta+\sin\theta)^{2}\cos\theta<0\text{ and }\mathcal{P}(\theta=\frac{\pi}{2})=\frac{\pi}{2}(\nu-1)^{2}>0.$
So $\phi_{e}^{\prime}(\theta)$ is negative before the only critical point of $\phi_{e}$ and positive after for $\theta\in(\arcsin\mu,\frac{\pi}{2})$. The only critical point of $\phi_{e}$ is a local minimum. That shows $\phi_{e}$ for $\theta\in(\arcsin\mu,\frac{\pi}{2})$ has no local maximum. ∎